### v3

# Engineering Software 2.0 by Interpolating Neural Networks: Unifying Training, Solving, and Calibration

Abstract
The evolution of artificial intelligence (AI) and neural network theories has revolutionized the way software is programmed, shifting from a hard-coded series of codes, Software 1.0, to a vast neural network, Software 2.0. However, this transition in engineering software has faced challenges such as data scarcity, multi-modality of data, low model accuracy, and slow inference. Here, we propose a new network based on interpolation theories and tensor decomposition, the interpolating neural network (INN) to open the new era of Engineering Software 2.0 that unifies training, solving, and calibration. Instead of interpolating training data, a common notion in computer science, INN interpolates grid points in the physical space whose coordinates and values are trainable. INN features orders of magnitude fewer trainable parameters (or degrees of freedom for solving), faster training/solving, less inference cost, smaller memory footprint, and higher model accuracy compared to multi-layer perceptron (MLP) or physics-informed neural networks (PINN). Various numerical experiments that cover computer science and engineering domains demonstrate that INN can solve over Zetta scale $(10^{21})$ partial differential equations and train/calibrate a dataset with extraordinary accuracy but fewer parameters using only a single graphics processing unit (GPU).
1 Introduction
The evolution of software programming methodologies has transitioned from reliance on rigidly hard-coded instructions, governed by explicitly defined human rules, to the adoption of neural network-based algorithms. The transition is coined as "from Software 1.0 to Software 2.0" [1]. The shift towards Software 2.0 has facilitated significant advancements in the domain of large language models (LLMs) and other foundational models [2, 3]. However, the application of these technologies within the field of engineering and science introduces distinct challenges. These challenges include scarcity and multi-modal aspect of data, low generalization error compared to classical numerical methods [4, 5, 6, 7], slow inference (function evaluation) time that challenges real-time control [8, 9, 10, 11] and statistical modeling, and resource-heavy model architecture [12, 13, 14, 15].
In this article, we propose the concept of Engineering Software 2.0, which aims to advance the current generation of computational and data-driven engineering software into a new paradigm.
“Engineering Software 2.0 is an end-to-end software system that unifies training, solving, and calibration in science/engineering problems."
As one promising enabler of Engineering Software 2.0, we introduce interpolating neural networks (INN). The key ideas behind INN are: 1) discretize an input domain into non-overlapping segments whose bounds are denoted by interpolation nodes, 2) construct a graph with the interpolation nodes and formulate the message-passing operation as a form of classical interpolation functions, and 3) optimize the values and coordinates of the interpolation nodes. Figure 1 shows the pictorial explanation of the approach. Note that one can also extrapolate for a certain range if the interpolation points are located outside of the range of training data. INN facilitates three major tasks in computational science and engineering: 1) train (or learn) on both experimental and simulation data with very high accuracy and minimal memory footprint, 2) solve high-dimensional partial differential equations (PDEs) with statistics involved, and 3) calibrate (or solve inverse problems) for uncertain parameters.
When the input domain is discretized with a regular mesh (see the special case of Figure 1), INN leverages the separation of variables and tensor decomposition (TD) [16, 17, 18, 19] to convert the growth rate of the computational cost of a high-dimensional problem from exponential to linear. This enables solving extremely challenging (over exascale, $10^{18}$) high-dimensional PDEs or training data with hundreds of input variables within the interpolation framework.
Our results demonstrate that INN can markedly enhance training/inference speed and model accuracy compared to the multilayer perceptron (MLP) or physics-informed neural network (PINN)[20] by dramatically reducing the number of trainable parameters and network connectivity. This also results in a huge reduction in the usage of random-access memory (RAM) and disk storage, facilitating the transferability of a trained/solved model across multiple workstations. Furthermore, INN can adapt its network structure further to improve model accuracy or incorporate statistical behavior. INN has a huge potential of replacing MLP and convolutional neural network (CNN) [21] in modern AI models with the aforementioned benefits.
Figure 1: Overview of Engineering Software 2.0 enabled with Interpolating Neural Networks (INN). The INN box illustrates the graph representation of the input domain discretized with an arbitrary Euclidean mesh (left) and a regular mesh as a special case (right). Node and edge features are given from the discretized mesh. After $Q$-step message passing, each node $j$ will store an interpolation function $N^{[Q](j)}(\bm{x})$. Finally, the readout operation sums the product of the interpolation functions and nodal values $\bm{u}^{(j)}$. The superscripts with square brackets [] and parentheses () denote the message passing step and graph node index, respectively. The interpolation operator $\mathcal{J}$ denotes that $\mathcal{J}\bm{u}(\bm{x})$ is a function that interpolates discrete values of $\bm{u}^{(j)}$. The INN trainer employs data-driven loss functions (e.g., mean squared error loss for regression) while the INN solver adopts a residual loss of a partial differential equation (PDE). A trained/solved INN model can then be employed as a forward model of a calibrator to solve an inverse problem.
2 INN: Unifying Training/Learning, Solving, and Calibration
Interpolating Neural Network (INN) generalizes numerical methods such as finite element method (FEM) [22, 23, 24] and neural networks such as multi-layer perception (MLP) as a graph network operation. Although the two (i.e., numerical methods and neural networks) have been developed by seemingly different groups of researchers, they share one thing in common; the goal is to find a function. The former finds a solution field of a differential equation while the latter finds an input-output relationship. INN is built upon this similarity and forms the backbone of Engineering Software 2.0 that unifies training, solving, and calibration, as illustrated in Figure 1.
Consider a regression problem that relates $I$ inputs and $L$ outputs. The first step of INN is to discretize the input domain in the $I$-dimensional Euclidean space into a mesh, which can be as general as an unstructured irregular mesh or as specific as a structured regular mesh. In any case, a mesh in the Euclidean space can be readily represented as a graph – the most general form of a discretized input, as illustrated in the INN box of Figure 1. Note that this conversion from mesh to graph has been widely accepted in literature [25, 26, 27]. Suppose there are $J$ nodes (or vertices) and $E$ edges in the graph, where the input domain is discretized with $C$ non-overlapping segments. Each node (e.g., node $j$) has two features: nodal coordinates $(\bm{x}^{(j)}\in\mathbb{R}^{I})$ and values $(\bm{u}^{(j)}\in\mathbb{R}^{L})$. Note that the superscript with parentheses refers to the graph node index. An edge that links node $j$ and $k$ has a feature $\bm{e}^{(j,k)}$, which stores indices of segments connected to the edge. For example, when $I=3,L=2$, node $10$ in Figure 1 stores $\bm{x}^{(10)}=(x^{(10)}_{1},x^{(10)}_{2},x^{(10)}_{3})$ and $\bm{u}^{(10)}=(u^{(10)}_{1},u^{(10)}_{2})$. The edge connecting nodes $30$ and $12$ stores $\bm{e}^{(30,12)}=\{5,7\}$.
A typical message passing in a graph neural network (GNN) returns a hidden state for each graph node which is mostly a tensor (including matrix and vector) [28]. In contrast, INN message passing returns a function (i.e., interpolation function $N^{(j)}(\bm{x})$, or often called shape function in computational mechanics) for each node $(j=1,\dots,J)$ as a hidden state. In other words, the message passing operation of INN is functional, an operation that returns a function. A general $Q$-step message passing of INN, $\mathcal{M}^{[q]}(*)$, can be expressed as:
$N^{[q](j)}(\bm{x})=\mathcal{M}^{[q]}\left(\bm{x}^{(j)},\{\bm{x}^{(k)},\bm{e}^{(j,k)}|k\in\mathcal{N}^{s}(j)\},N^{[q-1](j)}(\bm{x})\right),\quad q=1,\cdots,Q,\quad N^{[0](j)}(\bm{x})=0,$ (1)
where $N^{[q](j)}(\bm{x})$ is the interpolation function at node $j$ after $q$-th message passing and $\mathcal{N}^{s}(j)$ is a set of neighboring nodes of the center node $j$ with $s$ connections (i.e., $s$-hops, see Supplementary Information for visual illustrations). The operation $\mathcal{M}^{[q]}(*)$ constructs an interpolation function $N^{[q](j)}(\bm{x})$ for a graph node $\bm{x}^{(j)}$ using neighboring nodal coordinates and edge information: ${\bm{x}^{(k)},\bm{e}^{(j,k)}}$. This operation can be any interpolation method such as Lagrange polynomials [29, 22], meshfree [30, 31, 32], and splines [33, 34]. The $Q$-step message passing implies that it is also possible to progressively construct nonlinear or more problem-specific interpolation functions, theoretically based on the generalized finite element method [35, 36] and convolution hierarchical deep learning neural networks (C-HiDeNN) [7, 18]. Depending on the choice of interpolation technique, other hyper-parameters can be involved in $\mathcal{M}^{[q]}(*)$ such as the activation (or basis) function and dilation parameter [7, 18]. See Supplementary Information for various choices of the message passing operation.
During the forward propagation, an input variable $\bm{x}\in\mathbb{R}^{I}$ enters each graph node’s interpolation function $N^{[Q](j)}(\bm{x})$, followed by a graph-level readout operation:
$\mathcal{J}\bm{u}(\bm{x})=\sum^{J}_{j=1}{N^{(j)}(\bm{x})\bm{u}^{(j)}},\quad\bm{x}\in\mathbb{R}^{I},\bm{u}\in\mathbb{R}^{L}.$ (2)
Here, the superscript $[Q]$ is dropped for brevity (i.e., $N^{[Q](j)}(\bm{x})=N^{(j)}(\bm{x})$). The interpolation operator $\mathcal{J}$ will designate an interpolated field output throughout this paper. The readout operation can be written as a tensor contraction (or matrix multiplication):
$\mathcal{J}\bm{u}(\bm{x})=\begin{bmatrix}|&|&|&|\\ \bm{u}^{(1)}&\bm{u}^{(2)}&\cdots&\bm{u}^{(J)}\\ |&|&|&|\end{bmatrix}\cdot\begin{bmatrix}N^{(1)}(\bm{x})\\ N^{(2)}(\bm{x})\\ \vdots\\ N^{(J)}(\bm{x})\end{bmatrix}=\bm{U}^{T}\times^{1}_{2}\bm{\mathcal{X}(x)}=\bm{U}^{T}\bm{\mathcal{X}(x)},$ (3)
where $\bm{u}^{(j)}\in\mathbb{R}^{L},\bm{U}\in\mathbb{R}^{J\times L},\bm{\mathcal{X}(x)}\in\mathbb{R}^{J}$. The $\bm{U}^{T}\in\mathbb{R}^{L\times J}$ refers to the matrix transpose. Here, $\bm{U}^{T}\times^{1}_{2}\bm{\mathcal{X}(x)}$ denotes the tensor contraction operation between the second order of tensor $\bm{U}^{T}$ and the first order of tensor $\bm{\mathcal{X}(x)}$, while $\bm{U}^{T}\bm{\mathcal{X}(x)}$ is a matrix-vector multiplication. We note that $\bm{U}$ is a stack of nodal values $\bm{u}^{j}$, while $\bm{\mathcal{X}(x)}$ is a function of $\bm{x}$ that is parameterized with nodal coordinates $\bm{x}^{(j)}$ in the message passing operation.
The graph node features (i.e., coordinate ($\bm{x}^{(j)}$) and value ($\bm{u}^{(j)}$)) are trainable parameters of INN. If the nodal coordinates ($\bm{x}^{(j)}$) are fixed, one can find nodal values ($\bm{u}^{(j)}$) without changing the discretization of the input domain. If the nodal coordinates are also updated, the optimization will adjust the domain discretization similar to the r-adaptivity in FEM [23]. Once the forward propagation is defined, the loss function is chosen based on the problem type: training, solving, and calibration.
Special case 1: regular mesh with Tucker decomposition
When the input domain is discretized with a regular mesh, we can significantly reduce the trainable parameters (or degrees of freedom, DoFs) by leveraging Tensor Decomposition (TD) [16]. One of the widely used TD methods is the Tucker decomposition [37, 38]. Consider a three-input ($I=3$) and one output ($L=1$) system and assume the input domain is discretized with $J=J_{1}\times J_{2}\times J_{3}$ nodes, as shown in the special case box of Figure 1. To facilitate tensor notation, the nodal values $\bm{u}$ will be denoted with left/right super/sub scripts, $\prescript{m}{i}{u}^{(j)}_{l}$, where $i\in\mathbb{N}^{I}$ is the input index, $m\in\mathbb{N}^{M_{i}}$ is the mode index, $l\in\mathbb{N}^{L}$ is the output index and $j\in\mathbb{N}^{J_{i}}$ is the nodal index.
The interpolated field $\mathcal{J}u(\bm{x})\in\mathbb{R}^{L=1}$ can be represented as a Tucker product:
$\mathcal{J}u(\bm{x})=\llbracket\bm{\mathcal{G}};(\mathcal{J}\prescript{}{1}{\bm{u}})^{T},(\mathcal{J}\prescript{}{2}{\bm{u}})^{T},(\mathcal{J}\prescript{}{3}{\bm{u}})^{T}\rrbracket=\bm{\mathcal{G}}\times^{2}_{1}(\mathcal{J}\prescript{}{1}{\bm{u}})^{T}\times^{2}_{2}(\mathcal{J}\prescript{}{2}{\bm{u}})^{T}\times^{2}_{3}(\mathcal{J}\prescript{}{3}{\bm{u}})^{T},$ (4)
where the core tensor $\bm{\mathcal{G}}\in\mathbb{R}^{M_{1}\times M_{2}\times M_{3}}$ is a full trainable matrix typically smaller than the original nodal discretization ($M_{1}<J_{1},M_{2}<J_{2},M_{3}<J_{3}$). This enables data compression. The $\mathcal{J}\prescript{}{i}{\bm{u}}(x_{i})\in\mathbb{R}^{M_{i}\times 1}$ is the 1-dimensional (1D) interpolated output of $i$-th input dimension over $M_{i}$ modes represented as:
$\mathcal{J}\prescript{}{i}{\bm{u}}(x_{i})=\begin{bmatrix}\prescript{1}{i}{\bm{u}}^{(1)}&\prescript{1}{i}{\bm{u}}^{(2)}&\cdots&\prescript{1}{i}{\bm{u}}^{(J_{i})}\\ \prescript{2}{i}{\bm{u}}^{(1)}&\prescript{2}{i}{\bm{u}}^{(2)}&\cdots&\prescript{2}{i}{\bm{u}}^{(J_{i})}\\ \vdots&\vdots&\vdots&\vdots\\ \prescript{M_{i}}{i}{\bm{u}}^{(1)}&\prescript{M_{i}}{i}{\bm{u}}^{(2)}&\cdots&\prescript{M_{i}}{i}{\bm{u}}^{(J_{i})}\end{bmatrix}\cdot\begin{bmatrix}N^{(1)}(x_{i})\\ N^{(2)}(x_{i})\\ \vdots\\ N^{(J_{i})}(x_{i})\end{bmatrix}=\prescript{}{i}{\bm{U}}\times^{1}_{2}\prescript{}{i}{\bm{\mathcal{X}(x)}}=\prescript{}{i}{\bm{U}}\prescript{}{i}{\bm{\mathcal{X}(x)}},$ (5)
where $\prescript{}{i}{\bm{U}}\in\mathbb{R}^{M_{i}\times J_{i}}$ and $\prescript{}{i}{\bm{\mathcal{X}(x)}}\in\mathbb{R}^{J_{i}\times 1}$. When there is more than one output $(L>1)$, the interpolated field becomes a vector of $L$ elements:
$\mathcal{J}\bm{u}(\bm{x})$ $=\left[\mathcal{J}u_{1}(\bm{x}),\mathcal{J}u_{2}(\bm{x}),\cdots\mathcal{J}u_{L}(\bm{x})\right],\textit{where}$ (6)
$\mathcal{J}u_{l}(\bm{x})$ $=\llbracket\bm{\mathcal{G}}_{l};(\mathcal{J}\prescript{}{1}{\bm{u}}_{l})^{T},(\mathcal{J}\prescript{}{2}{\bm{u}}_{l})^{T},(\mathcal{J}\prescript{}{3}{\bm{u}}_{l})^{T}\rrbracket=\bm{\mathcal{G}}_{l}\times^{2}_{1}(\mathcal{J}\prescript{}{1}{\bm{u}}_{l})^{T}\times^{2}_{2}(\mathcal{J}\prescript{}{2}{\bm{u}}_{l})^{T}\times^{2}_{3}(\mathcal{J}\prescript{}{3}{\bm{u}}_{l})^{T}.$
Special case 2: regular mesh with CP decomposition
Tucker decomposition can be further simplified to CANDECOMP/PARAFAC (CP) decomposition [39, 40, 41] by setting the core tensor $\bm{\mathcal{G}}$ as an order-$I$ super diagonal tensor: $\bm{\mathcal{G}}\in\mathbb{R}^{M^{I}}$, $M=M_{1}=\cdots=M_{I}$, all zero entries except the diagonal elements. If we further set the diagonal elements of $\bm{\mathcal{G}}$ to be 1, the CP decomposition yields:
$\mathcal{J}\bm{u}(\bm{x})$ $=\llbracket\bm{\mathcal{G}};(\mathcal{J}\prescript{}{1}{\bm{u}})^{T},(\mathcal{J}\prescript{}{2}{\bm{u}})^{T},(\mathcal{J}\prescript{}{3}{\bm{u}})^{T}\rrbracket=\sum^{M}_{m=1}{\prod^{I}_{i=1}{\mathcal{J}\prescript{m}{i}{u}(x_{i})}},$ (7)
an inner product of the three interpolated vectors of length $M$.
It is important to note that the interpolation function $N^{(j)}(x_{i})$ in Eq.5 is a one-dimensional function. In other words, both Tucker and CP decomposition replace a high-dimensional interpolation with multiple one-dimensional interpolations that dramatically reduce the number of trainable parameters (or DoFs). As given in Table 1, the number of trainable parameters of the full interpolation scales exponentially with the discretization $J_{i}$, whereas the Tucker and the CP decomposition scale linearly with $J_{i}$.
Table 1: Number of trainable parameters (or degrees of freedom, DoFs) of full interpolation, Tucker decomposition, and CP decomposition for an $I$-input single output relationship. The input domain is discretized with a regular grid of $J_{1}\times J_{2}\times\cdots\times J_{I}$ discretization. Assume that the nodal coordinates $(\bm{x}^{j})$ are fixed.
Full interpolation Tucker decomposition CP decomposition
Trainable parameters $\prod^{I}_{i}{J_{i}}$ $\prod^{I}_{i}{M_{i}}+\sum^{I}_{i}{M_{i}J_{i}}$ $M\sum^{I}_{i}{J_{i}}$
INN trainer
A general INN forward propagation is provided in Eq.3 as a tensor contraction. The INN trainer optimizes nodal values $\bm{U}=\{\bm{u}^{(j)}\}_{j=1,\cdots,J}$ (and if needed, nodal coordinates ($\bm{X}=\{\bm{x}^{(j)}\}_{j=1,\cdots,J}$)) under a given loss function and training data. Consider a regression problem with $K$ labeled data: $(\bm{x}^{*}_{k},\bm{u}^{*}_{k})$, $k=1,\cdots,K$ and $\bm{x}^{*}_{k}\in\mathbb{R}^{I},\bm{u}^{*}_{k}\in\mathbb{R}^{L}$. Here the superscript * denotes data. A mean squared error (MSE) loss function for this regression problem is defined as:
$loss(\bm{U},\bm{X})=\frac{1}{K}\sum_{k}(\mathcal{J}\bm{u}(\bm{x}^{*}_{k})-\bm{u}^{*}_{k})^{2}.$ (8)
Finally, an optimization problem is formulated as:
$\underset{\bm{U},\bm{X}}{\text{minimize}}\ loss(\bm{U},\bm{X}).$ (9)
Several benchmark problems are introduced in the Supplementary Information. The superior training accuracy, faster training and inference time, and fewer trainable parameters of the INN trainer are demonstrated.
INN solver
INN solver generalizes classical numerical methods such as FEM and meshfree, as well as the model order reduction methods such as proper generalized decomposition (PGD) [42] and tensor decomposition (TD) [17, 43]. Here, we introduce a formulation with CP decomposition to solve the generalized space $(\bm{x})$ time $(t)$ parameter $(\bm{\theta})$ PDE, which is prohibitive for most numerical methods and machine learning approaches [42]. Consider a parameterized space-time PDE:
$\mathcal{L}u(\bm{x},t,\bm{\theta})=0,$ (10)
where $\mathcal{L}$ is the general partial differential operator (can be linear or nonlinear).
The INN solver uses the same neural network structure as the trainer, but the definition of loss function differs since training data is not known before solving the equation. Instead, the loss function is formulated as the weighted summation of the PDE residual [44, 45, 46]. INN approximates the general space $(\bm{x})$ time $(t)$ parameter $(\bm{\theta})$ solution using the CP decomposition format:
$\begin{split}\mathcal{J}{\bm{u}}(\bm{x},t,\bm{\theta};\bm{U},\bm{X})&=\sum_{m=1}^{M}[\mathcal{J}\prescript{m}{x_{1}}{\bm{u}}(x_{1})]\odot\cdots\odot[\mathcal{J}\prescript{m}{x_{d}}{\bm{u}}(x_{d})]\odot[\mathcal{J}\prescript{m}{t}{\bm{u}}(x_{t})]\\ &\odot[\mathcal{J}\prescript{m}{\theta_{1}}{\bm{u}}(\theta_{1})]\odot\cdots\odot[\mathcal{J}\prescript{m}{\theta_{k}}{\bm{u}}(\theta_{k})].\end{split}$ (11)
where $d$ and $k$ are the spatial dimension and the number of parameters, respectively. The operator $\odot$ refers to the element-wise multiplication. Similar to the trainer, the goal is to find nodal values $\bm{U}$ (and nodal coordinates $\bm{X}$, if one wants to adapt the mesh). As a result, INN obtains the space-time-parameter solution by minimizing the loss function:
[…]
where $\delta\bm{u}$ is the weight function and can be defined using Galerkin, Petrov-Galerkin, collocation or quadratic formulation. It should be noted that INN should automatically satisfy Dirichlet and initial conditions. As a result, only weighted sum residual residual is considered in the loss function Eq. 11. See the Supplementary Information for detailed derivations of the loss function.
In Supplementary Information, a 3D linear elasticity problem is solved with the full interpolation INN and CP decomposition INN, and compared with FEM. Especially the INN with CP decomposition enables a voxel mesh of 10 million cube - $10^{21}$ grid points, which expands the state-of-the-art FEM limit by order of 10.
INN calibrator
[…]
$\bm{x}^{*}=\underset{\bm{x}}{\text{argmin}}\frac{1}{K}\sum_{k}(\mathcal{J}\bm{u}(\bm{x})-\ \bm{u}^{*}_{k})^{2},$ (13)
where $\mathcal{J}\bm{u}(\bm{x})$ is the trained or solved forward model, $\bm{u}^{*}_{k}$ is the $k$-th measured observation and $\bm{x}^{*}$ is the calibrated input variable.
A few calibration benchmarks are provided in Supplementary Information that demonstrate the superior calibration accuracy and reliability of the INN calibrator.
3 Scientific and Engineering applications
Figure 2: A summary of engineering applications of the INN methods to scientific and engineering problems. The calibration problem develops a reduced-order model of laser powder bed additive manufacturing to calibrate the heat source parameters from experimental data. A space-time-parameter solution of the same problem with large parameter space is demonstrated in Solving category. INN solvers result in significant storage reduction and fast simulation. Finally, a data-driven real-time online monitoring and feedback control tool is formulated with INN that uses only 18% training parameters and is 18-31 times faster compared to MLP.
This section will illustrate how Engineering Software 2.0 enabled with INN can be applied to model various facets of metal additive manufacturing by integrating training, solving, and calibration processes. Note that the INN with CP decomposition will be referred to INN for brevity. Laser powder bed fusion (L-PBF) is a type of metal additive manufacturing in which metal powders are layered and melted by a laser beam to create objects that match the provided CAD designs [47]. The added flexibility of L-PBF comes with a very large design space and spatially varying microstructure after manufacturing. Therefore, considerable research is being devoted to employing in situ monitoring data for real-time control of L-PBF processes to minimize the variability [8]. Computational modeling of L-PBF spans the simulation of the manufacturing process with a high dimensional design space (solving), computing the internal relationship among different manufacturing parameters (calibrating), and designing an online monitoring and control system for L-PBF (training).
The governing equation for modeling the manufacturing process is given by
$\frac{\partial T(\mathbf{x},t)}{\partial t}-k\Delta T(\mathbf{x},t)=Q_{h}(\mathbf{x},t,P,R,\eta,d)$ (14)
This is a heat conduction equation where the moving laser heat source is incorporated through the source term, $Q_{h}$. The equation involves spatial parameter $\mathbf{x}$, temporal variable $t$, thermal conductivity $k$, power $P$, beam radius $R$, absorptivity $\eta$, and laser depth $d$. For the simplest case of single track laser pass, the initial condition is $T(\mathbf{x},0)=T_{0}(\mathbf{x})$, and the boundary conditions (BC) are $T=\tilde{T}$ in $\partial\Omega_{T}$ for Dirichlet BC and $\frac{\partial T}{\partial\mathbf{x}}\cdot\mathbf{n}=\bar{q}$ in $\partial\Omega_{q}$ for Neumann BC, where the spatial domain boundary $\partial\Omega=\partial\Omega_{T}\cup\partial\Omega_{q}$.
3.1 Calibration of Heat Source Parameters for Additive Manufacturing
In this example, the INN is used for calibrating a model of heat source of L-PBF, which uses a laser heat source to selectively melt metal powder bed to form a connected shape. A commonly used Gaussian beam profile can be modeled as a volumetric heat source [48]. The heat source equation reads,
$Q_{h}(\textbf{x},t,\textbf{p})=\frac{2\mathcal{P}\eta}{\pi r^{2}_{b}d}e^{-\frac{2(x^{2}+y^{2})}{r^{2}_{b}}}$ (15)
where, $z_{top}-z\leq d$, $z_{top}$ is z coordinate of the current powder bed surface. The notations follow Eq. 14 and $r_{b}$ is the standard deviation of the Gaussian source. To get an accurate representation of the physics behind it, model parameters need to be calibrated against experimental data. Following [49], three calibration parameters $p_{1}$, $p_{2}$, and $p_{3}$ can be defined as, $d=p_{1}\frac{\mathcal{P}}{V}RHF^{2}$, $\eta=p_{2}\frac{\mathcal{P}}{V}RHF^{2}$, and $r_{b}=p_{3}\frac{\mathcal{P}}{V}RHF^{2}$. Here, $P$ is laser power, $V$ is the scan speed, and $RHF$ is a residual heat measure predetermined based on a laser toolpath. The challenges for calibrating these parameters are: 1) the computational models are often very expensive to perform even one forward run, 2) the computational models are not automatically differentiable, and 3) limited experimental data is available. As a demonstration for this problem, INN with 1760 trainable parameters is trained on 44 sets of meltpool temperature data generated by the finite difference method (FDM) using 4 different power-scanning speed setup. Later, this trained INN is used for optimizing the three aforementioned parameters using gradient-based optimizers. To show the superiority of INN over traditional neural networks in such data-scarce applications, a 3-layer MLP with 2689 trainable parameters is trained using the same data fed to INN. Details regarding this calibration process are provided in the Supplementary Information.
Both INN-based and MLP-based calibrations are repeated 50 times to give a statistical comparison (see Figure 2). Even with only 44 sets of simulation data, the INN calibrated parameters that produce mean meltpool temperature (or TEP, thermal emission plank) within 7.9% of the experimental data. On the contrary, MLP produced a difference of 44.1% against experimental data, which shows that it hardly learns the underlying physics using such small data with even more trainable parameters.
3.2 Solving High Dimensional Space-Time-Parameter Heat Equation
The governing equation Eq.14 has 8 input variables; 3 for spatial domain ($x,y,z$), 1 for time domain ($t$), and 4 for parametric domain ($k,p,e,d$). See Supplementary Information for a detailed problem definition. This equation is not feasible for process design with most existing computational methods, including the data-driven method inspired by the finite element method (FEM) and the physics-informed neural network (PINN) [20].
To solve the problem with FEM-based data-driven method, one first needs to sample from the 4-dimensional parametric domain, followed by running the time-stepping-based finite element analysis for each sampled parameter set. Then the FEM solutions should be stored as offline training data, and a neural network surrogate model is trained using the data. The computational cost for such operations is prohibitive since large space-time mesh is required in each simulation for real engineering application, as detailed in the Supplementary Information. Moreover, as a larger number of training data is generated due to the high dimensionality of the input domain, this method becomes impractical in terms of data storage. Furthermore, running a large-scale FEM simulation requires a large amount of computational memory.
PINN is a unified framework of training and solving for such parameterized PDEs; however, the computational cost still grows exponentially as the number of input variables grows. As illustrated in Supplementary Information, it is very challenging to solve this 8-dimensional equation. In addition, the convergence of solution accuracy is not guaranteed especially for such high-dimensional problems. Thus, it remains to be seen if PINN can even find a good solution for more than 7 input variables. PINN also suffers from memory issues for GPU training due to the exponential growth of the collocation points although batch training can be adopted to alleviate memory usage.
Contrary to FEM and PINN, INN solves this space-time-parameter problem all-at-once. As it is based on the projected 1D interpolation, the number of trainable parameters increases linearly with the input variables, making the computational cost manageable. Meanwhile, since tensor-decomposed data structures are leveraged, INN has the benefits of significant memory efficiency and storage gain, especially for high-dimensional problems. In Figure 2, we solve the 8-dimensional space-time-parameter PDE with the total number of space-time-parameter grid points of $3\times 10^{32}$, where we use 25,000 grid points to discretize each spatial and parametric coordinate and 50 grid points for the temporal coordinate. The details of the solver example are provided in the Supplementary Information. Since the number of trainable parameters of INN grows linearly, INN has a significant storage gain of $1.7\times 10^{25}$ compared to the classical FEM.
Moreover, when using INN, once the parametric solution of the AM problem is obtained, one can easily investigate the uncertainty propagation by imposing probabilistic distributions on the inputs of INN, thanks to the tensor-decomposed data structure. Details on the statistical prediction of parametric solution can be found in the Supplementary Information.
3.3 Real-time Online Control of Laser Powder Bed Fusion Additive Manufacturing
In this application, the INN trainer is used to devise an online control system for L-PBF additive manufacturing. The goal is to maintain a homogeneous melt pool temperature across each layer. The hypothesis is that a homogeneous melt pool temperature across each layer of material deposition will reduce the variability in the microstructure and mechanical properties. The fundamental challenge of implementing a model predictive control system is the computational time required for the forward prediction (involving finite element or computational fluid dynamics models). For prediction, the data is generated by fusing experiments and computational methods (see Supplementary Information for details). The online control loop uses the trained INN model to inform the manufacturing machine. There are two models: forward model and inverse model.
The forward model uses the mechanistic features (precomputed based on the tool path and geometry of the build to generalize a data-driven model for unseen tool path and geometric features; more details can be found in [50]), and melt pool temperature of the previous layer (measured by the sensor as Thermal Emission Plank (TEP) signal; more details in the Supplementary Information) to predict the TEP/melt pool temperature in the next layer for an uncontrolled process. The uncontrolled TEP is converted to a homogenized TEP over the entire layer. The inverse model takes this targeted, homogenized TEP as the input along with the previous layer’s mechanistic features to predict the required laser power for the next layer. The laser power is then exported to the machine to print the next layer. The forward model is trained on experimentally observed data while the inverse model is trained on computational data coming from a finite element solver. The INN model is applied for both forward and inverse prediction and compared against MLP. Figure 2 shows that for the inverse model INN is at least 18 to 31 times faster to train compared to the MLP to reach the same level of training error, and it uses only about 18% of the parameters that MLP uses. The speed-up depends on the amount of data used. The INN method is consistently more accurate compared to MLP (see Supplementary Information for more details).
4 Conclusions
We show that interpolating neural networks (INN) can train, solve, and calibrate scientific and engineering problems that are extremely challenging or prohibitive for existing numerical methods or machine learning models. The keys to INN’s success in those problems are 1) it interpolates virtual points with well-established interpolation theories (not interpolating training data with a black-box network) and 2) it leverages the idea of separation of variables and tensor decomposition to resolve the curse of dimensionality. Due to the significantly reduced trainable parameters without losing approximation capabilities, INN can train/solve/calibrate faster with less memory requirement, and achieve extremely high model accuracy and fast inference time.
As this is the original article on INN, we have myriads of research questions that may gather huge research interests across various areas including but not limited to scientific machine learning, applied mathematics, computer vision, and data science. To mention a few:
  • Multi-resolution aspect (i.e., a varying number of segments across INN networks, $M$) of INN needs to be investigated. See Supplementary Information for further discussions.
  • Convergence of INN for training can be proved mathematically.
  • INN training on ensembles of training data might be a way to enhance the generality of the model for image problems or noisy data.
  • Since INN could be used as both a solver for physics-based problems and a function approximator for data-driven problems, it is of interest to combine the two behaviors together into one single model, so that we can solve large scale problems involving both complex physics and scarce data. For example, the calibration task of LPBF heat source.
We expect INN can open a next-generation Engineering Software 2.0 by theoretically integrating training, solving, and calibration and tackling extremely challenging problems in the science and engineering domain.

### v4

@@
-Engineering Software 2.0 by Interpolating Neural Networks: Unifying Training, Solving, and Calibration
+Interpolating neural network: A lightweight yet precise architecture for data training, equation solving, and parameter calibration
 Abstract
-The evolution of artificial intelligence (AI) and neural network theories has revolutionized the way software is programmed, shifting from a hard-coded series of codes, Software 1.0, to a vast neural network, Software 2.0.
-However, this transition in engineering software has faced challenges such as data scarcity, multi-modality of data, low model accuracy, and slow inference.
-Here, we propose a new network based on interpolation theories and tensor decomposition, the interpolating neural network (INN) to open the new era of Engineering Software 2.0 that unifies training, solving, and calibration.
-Instead of interpolating training data, a common notion in computer science, INN interpolates grid points in the physical space whose coordinates and values are trainable.
-INN features orders of magnitude fewer trainable parameters (or degrees of freedom for solving), faster training/solving, less inference cost, smaller memory footprint, and higher model accuracy compared to multi-layer perceptron (MLP) or physics-informed neural networks (PINN).
-Various numerical experiments that cover computer science and engineering domains demonstrate that INN can solve over Zetta scale $(10^{21})$ partial differential equations and train/calibrate a dataset with extraordinary accuracy but fewer parameters using only a single graphics processing unit (GPU).
+Artificial intelligence (AI) has revolutionized software development, shifting from task-specific codes (Software 1.0) to neural network-based approaches (Software 2.0).
+However, applying this transition in engineering software presents challenges, including low surrogate model accuracy, the curse of dimensionality in inverse design, and rising complexity in physical simulations.
+We introduce an interpolating neural network (INN), grounded in interpolation theory and tensor decomposition, to realize Engineering Software 2.0 by advancing data training, partial differential equation solving, and parameter calibration.
+INN offers orders of magnitude fewer trainable/solvable parameters for comparable model accuracy than traditional multi-layer perceptron (MLP) or physics-informed neural networks (PINN).
+Demonstrated in metal additive manufacturing, INN rapidly constructs an accurate surrogate model of Laser Powder Bed Fusion (L-PBF) heat transfer simulation, achieving sub-10-micrometer resolution for a 10 mm path in under 15 minutes on a single GPU.
+This makes a transformative step forward across all domains essential to engineering software.
 1 Introduction
@@
 The transition is coined as "from Software 1.0 to Software 2.0" [1].
-The shift towards Software 2.0 has facilitated significant advancements in the domain of large language models (LLMs) and other foundational models [2, 3].
-However, the application of these technologies within the field of engineering and science introduces distinct challenges.
-These challenges include scarcity and multi-modal aspect of data, low generalization error compared to classical numerical methods [4, 5, 6, 7], slow inference (function evaluation) time that challenges real-time control [8, 9, 10, 11] and statistical modeling, and resource-heavy model architecture [12, 13, 14, 15].
+The shift towards Software 2.0 partially resolved the issue of labor-intensive programming in Software 1.0 and has significantly advanced the domain of large language models (LLMs) and other foundational models [2, 3].
+However, the application of these technologies in the fields of engineering and science presents unique challenges.
+One such challenge is the relatively low generalization accuracy of machine learning (ML)-based solvers compared to traditional numerical methods [4, 5, 6, 7, 8].
+After in-depth research was conducted on those methods, an increasing number of researchers began to question their effectiveness [8, 9].
+Another challenge is the curse of dimensionality that arises when solving inverse design problems.
+As the design parameter space grows, computation requirements easily become prohibitive.
+Furthermore, the complexity of physical simulations continues to increase.
+For instance, additive manufacturing (AM) or integrated circuit (IC) simulations necessitate extremely fine resolution with high accuracy and multi-physics/multi-scale analysis within a reasonable simulation time [10, 11, 12].
 In this article, we propose the concept of Engineering Software 2.0, which aims to advance the current generation of computational and data-driven engineering software into a new paradigm.
-“Engineering Software 2.0 is an end-to-end software system that unifies training, solving, and calibration in science/engineering problems."
-As one promising enabler of Engineering Software 2.0, we introduce interpolating neural networks (INN).
-The key ideas behind INN are: 1) discretize an input domain into non-overlapping segments whose bounds are denoted by interpolation nodes, 2) construct a graph with the interpolation nodes and formulate the message-passing operation as a form of classical interpolation functions, and 3) optimize the values and coordinates of the interpolation nodes.
-Figure 1 shows the pictorial explanation of the approach.
-Note that one can also extrapolate for a certain range if the interpolation points are located outside of the range of training data.
-INN facilitates three major tasks in computational science and engineering: 1) train (or learn) on both experimental and simulation data with very high accuracy and minimal memory footprint, 2) solve high-dimensional partial differential equations (PDEs) with statistics involved, and 3) calibrate (or solve inverse problems) for uncertain parameters.
-When the input domain is discretized with a regular mesh (see the special case of Figure 1), INN leverages the separation of variables and tensor decomposition (TD) [16, 17, 18, 19] to convert the growth rate of the computational cost of a high-dimensional problem from exponential to linear.
-This enables solving extremely challenging (over exascale, $10^{18}$) high-dimensional PDEs or training data with hundreds of input variables within the interpolation framework.
-Our results demonstrate that INN can markedly enhance training/inference speed and model accuracy compared to the multilayer perceptron (MLP) or physics-informed neural network (PINN)[20] by dramatically reducing the number of trainable parameters and network connectivity.
-This also results in a huge reduction in the usage of random-access memory (RAM) and disk storage, facilitating the transferability of a trained/solved model across multiple workstations.
-Furthermore, INN can adapt its network structure further to improve model accuracy or incorporate statistical behavior.
-INN has a huge potential of replacing MLP and convolutional neural network (CNN) [21] in modern AI models with the aforementioned benefits.
-Figure 1: Overview of Engineering Software 2.0 enabled with Interpolating Neural Networks (INN).
+“Engineering Software 2.0 is an end-to-end software system that unifies training, solving, and calibrating in science/engineering problems."
+We introduce interpolating neural networks (INNs) to achieve the new paradigm of Engineering Software 2.0.
+The key ideas behind INNs are: 1) discretize an input domain into non-overlapping segments whose bounds are denoted by interpolation nodes, 2) construct a graph with the interpolation nodes and formulate the message passing operation as a form of interpolation functions, and 3) optimize the values and coordinates of the interpolation nodes for a given loss function.
+Figure 1 shows a pictorial description of the approach.
+When the input domain is discretized with a regular mesh (see the special case in Figure 1), an INN leverages tensor decomposition (TD) [13, 14, 15, 16] to convert the growth of the computational cost of a high-dimensional problem from exponential to linear.
+INN facilitates three major tasks in computational science and engineering: 1) trains on both experimental and simulation data with very high accuracy and minimal memory footprint, 2) solves high-dimensional parameterized partial differential equations (PDEs), and 3) calibrates uncertain parameters rapidly and accurately.
+[…]
+Figure 1: Overview of Engineering Software 2.0 enabled with Interpolating Neural Network (INN).
 The INN box illustrates the graph representation of the input domain discretized with an arbitrary Euclidean mesh (left) and a regular mesh as a special case (right).
@@
 A trained/solved INN model can then be employed as a forward model of a calibrator to solve an inverse problem.
-2 INN: Unifying Training/Learning, Solving, and Calibration
-Interpolating Neural Network (INN) generalizes numerical methods such as finite element method (FEM) [22, 23, 24] and neural networks such as multi-layer perception (MLP) as a graph network operation.
-Although the two (i.e., numerical methods and neural networks) have been developed by seemingly different groups of researchers, they share one thing in common; the goal is to find a function.
+Figure 2: Benchmark comparison of the single-scale finite difference method (FDM) solver [18], the variational multiscale FEM solver [11], and the INN solver.
+The data points for the first two methods with dashed marker edges are estimated, while that of INN with a solid marker edge is computed.
+A 3D-space, time, and 4D-parameter heat transfer equation is solved to model a 10 mm single-track laser powder bed fusion (L-PBF) metal AM.
+Detailed problem definition and explanation can be found in Section 3.
+All solvers are graphics processing units (GPU) optimized with the JAX library [19].
+2 INN: Unification of Training/Learning, Solving, and Calibrating
+Interpolating Neural Network (INN) generalizes numerical methods such as the finite element method (FEM) [20, 21, 22] as a subset of deep neural network (DNN) via graph network operations.
+Although different groups of researchers have developed the two (FEM and DNN), they share one thing in common; the goal is to find a function.
 The former finds a solution field of a differential equation while the latter finds an input-output relationship.
-INN is built upon this similarity and forms the backbone of Engineering Software 2.0 that unifies training, solving, and calibration, as illustrated in Figure 1.
+INN is built upon this similarity and forms the backbone of Engineering Software 2.0 that unifies training, solving, and calibrating, as illustrated in Figure 1.
 Consider a regression problem that relates $I$ inputs and $L$ outputs.
-The first step of INN is to discretize the input domain in the $I$-dimensional Euclidean space into a mesh, which can be as general as an unstructured irregular mesh or as specific as a structured regular mesh.
-In any case, a mesh in the Euclidean space can be readily represented as a graph – the most general form of a discretized input, as illustrated in the INN box of Figure 1.
-Note that this conversion from mesh to graph has been widely accepted in literature [25, 26, 27].
+The first step of an INN is to discretize the input domain in the $I$-dimensional Euclidean space into a mesh, which can be as general as an unstructured irregular mesh or as specific as a structured regular mesh.
+In any case, a mesh in the Euclidean space can be readily represented as a graph – the most general form of a discretized input – as illustrated in the INN box of Figure 1.
+The conversion from mesh to graph has been widely accepted in the literature [23, 24, 25, 26].
 Suppose there are $J$ nodes (or vertices) and $E$ edges in the graph, where the input domain is discretized with $C$ non-overlapping segments.
+Each segment occupies a subspace of the $I$-dimensional input domain.
+For instance in Figure 1, the Euclidean mesh consists of segments of hexahedrons and triangular pillars.
 Each node (e.g., node $j$) has two features: nodal coordinates $(\bm{x}^{(j)}\in\mathbb{R}^{I})$ and values $(\bm{u}^{(j)}\in\mathbb{R}^{L})$.
 Note that the superscript with parentheses refers to the graph node index.
-An edge that links node $j$ and $k$ has a feature $\bm{e}^{(j,k)}$, which stores indices of segments connected to the edge.
+An edge that links node $j$ and $k$ has a feature $\bm{e}^{(j,k)}$ that stores indices of segments connected to the edge.
 For example, when $I=3,L=2$, node $10$ in Figure 1 stores $\bm{x}^{(10)}=(x^{(10)}_{1},x^{(10)}_{2},x^{(10)}_{3})$ and $\bm{u}^{(10)}=(u^{(10)}_{1},u^{(10)}_{2})$.
@@
 A typical message passing in a graph neural network (GNN) returns a hidden state for each graph node which is mostly a tensor (including matrix and vector) [27].
-In contrast, INN message passing returns a function (i.e., interpolation function $N^{(j)}(\bm{x})$, or often called shape function in computational mechanics) for each node $(j=1,\dots,J)$ as a hidden state.
-In other words, the message passing operation of INN is functional, an operation that returns a function.
-A general $Q$-step message passing of INN, $\mathcal{M}^{[q]}(*)$, can be expressed as:
+In contrast, the INN message passing returns a function (i.e., interpolation function $N^{(j)}(\bm{x})$, or often called shape function in computational mechanics) for each node $(j=1,\dots,J)$ as a hidden state.
+A general $Q$-step message passing of an INN, $\mathcal{M}^{[q]}(*)$, can be expressed as:
 $N^{[q](j)}(\bm{x})=\mathcal{M}^{[q]}\left(\bm{x}^{(j)},\{\bm{x}^{(k)},\bm{e}^{(j,k)}|k\in\mathcal{N}^{s}(j)\},N^{[q-1](j)}(\bm{x})\right),\quad q=1,\cdots,Q,\quad N^{[0](j)}(\bm{x})=0,$ (1)
-where $N^{[q](j)}(\bm{x})$ is the interpolation function at node $j$ after $q$-th message passing and $\mathcal{N}^{s}(j)$ is a set of neighboring nodes of the center node $j$ with $s$ connections (i.e., $s$-hops, see Supplementary Information for visual illustrations).
-The operation $\mathcal{M}^{[q]}(*)$ constructs an interpolation function $N^{[q](j)}(\bm{x})$ for a graph node $\bm{x}^{(j)}$ using neighboring nodal coordinates and edge information: ${\bm{x}^{(k)},\bm{e}^{(j,k)}}$.
-This operation can be any interpolation method such as Lagrange polynomials [29, 22], meshfree [30, 31, 32], and splines [33, 34].
-The $Q$-step message passing implies that it is also possible to progressively construct nonlinear or more problem-specific interpolation functions, theoretically based on the generalized finite element method [35, 36] and convolution hierarchical deep learning neural networks (C-HiDeNN) [7, 18].
+where $N^{[q](j)}(\bm{x})$ is the interpolation function at node $j$ after $q$-th message passing and $\mathcal{N}^{s}(j)$ is a set of neighboring nodes of the center node $j$ with $s$ connections (i.e., $s$-hops, see Supplementary Information (SI) Section 1.1 for visual illustrations).
+It is worth noting that the interpolation functions satisfy the Kronecker delta property, i.e., $N^{[q](j)}(\bm{x}^{(k)})=\delta_{jk}$.
+The operation $\mathcal{M}^{[q]}(*)$ constructs an interpolation function $N^{[q](j)}(\bm{x})$ for a graph node $\bm{x}^{(j)}$ using neighboring nodal coordinates and edge information: ${\bm{x}^{(k)},\bm{e}^{(j,k)}}$, and the interpolation function of the previous message passing: $N^{[q-1](j)}(\bm{x})$.
+This can be any interpolation method such as Lagrange polynomials [28, 20], meshfree [29, 30, 31], and splines [32, 33].
+The $Q$-step message passing implies that it is also possible to progressively construct nonlinear or more problem-specific interpolation functions, theoretically based on the generalized finite element method [34, 35] and convolution hierarchical deep learning neural networks [7, 15].
 Depending on the choice of interpolation technique, other hyper-parameters can be involved in $\mathcal{M}^{[q]}(*)$ such as the activation (or basis) function and dilation parameter [7, 15].
-See Supplementary Information for various choices of the message passing operation.
+See SI Section 1 for various choices of the message passing operation.
 During the forward propagation, an input variable $\bm{x}\in\mathbb{R}^{I}$ enters each graph node’s interpolation function $N^{[Q](j)}(\bm{x})$, followed by a graph-level readout operation:
-$\mathcal{J}\bm{u}(\bm{x})=\sum^{J}_{j=1}{N^{(j)}(\bm{x})\bm{u}^{(j)}},\quad\bm{x}\in\mathbb{R}^{I},\bm{u}\in\mathbb{R}^{L}.$ (2)
-Here, the superscript $[Q]$ is dropped for brevity (i.e., $N^{[Q](j)}(\bm{x})=N^{(j)}(\bm{x})$).
+$\mathcal{J}\bm{u}(\bm{x})=\sum^{J}_{j=1}{N^{(j)}(\bm{x})\bm{u}^{(j)}},\quad\bm{x}\in\mathbb{R}^{I},\bm{u}\in\mathbb{R}^{L},$ (2)
+where the superscript $[Q]$ is dropped for brevity (i.e., $N^{[Q](j)}(\bm{x})=N^{(j)}(\bm{x})$).
 The interpolation operator $\mathcal{J}$ will designate an interpolated field output throughout this paper.
 The readout operation can be written as a tensor contraction (or matrix multiplication):
-$\mathcal{J}\bm{u}(\bm{x})=\begin{bmatrix}|&|&|&|\\ \bm{u}^{(1)}&\bm{u}^{(2)}&\cdots&\bm{u}^{(J)}\\ |&|&|&|\end{bmatrix}\cdot\begin{bmatrix}N^{(1)}(\bm{x})\\ N^{(2)}(\bm{x})\\ \vdots\\ N^{(J)}(\bm{x})\end{bmatrix}=\bm{U}^{T}\times^{1}_{2}\bm{\mathcal{X}(x)}=\bm{U}^{T}\bm{\mathcal{X}(x)},$ (3)
-where $\bm{u}^{(j)}\in\mathbb{R}^{L},\bm{U}\in\mathbb{R}^{J\times L},\bm{\mathcal{X}(x)}\in\mathbb{R}^{J}$.
-The $\bm{U}^{T}\in\mathbb{R}^{L\times J}$ refers to the matrix transpose.
-Here, $\bm{U}^{T}\times^{1}_{2}\bm{\mathcal{X}(x)}$ denotes the tensor contraction operation between the second order of tensor $\bm{U}^{T}$ and the first order of tensor $\bm{\mathcal{X}(x)}$, while $\bm{U}^{T}\bm{\mathcal{X}(x)}$ is a matrix-vector multiplication.
-We note that $\bm{U}$ is a stack of nodal values $\bm{u}^{j}$, while $\bm{\mathcal{X}(x)}$ is a function of $\bm{x}$ that is parameterized with nodal coordinates $\bm{x}^{(j)}$ in the message passing operation.
-The graph node features (i.e., coordinate ($\bm{x}^{(j)}$) and value ($\bm{u}^{(j)}$)) are trainable parameters of INN.
+$\mathcal{J}\bm{u}(\bm{x})=\begin{bmatrix}|&|&|&|\\ \bm{u}^{(1)}&\bm{u}^{(2)}&\cdots&\bm{u}^{(J)}\\ |&|&|&|\end{bmatrix}\cdot\begin{bmatrix}N^{(1)}(\bm{x})\\ N^{(2)}(\bm{x})\\ \vdots\\ N^{(J)}(\bm{x})\end{bmatrix}=\bm{U}\times^{1}_{2}\bm{\mathcal{X}(x)}=\bm{U}\bm{\mathcal{X}(x)},$ (3)
+where $\bm{u}^{(j)}\in\mathbb{R}^{L},\bm{U}\in\mathbb{R}^{L\times J},\bm{\mathcal{X}(x)}\in\mathbb{R}^{J}$.
+Here, $\bm{U}\times^{1}_{2}\bm{\mathcal{X}(x)}$ denotes the tensor contraction operation between the second dimension of tensor $\bm{U}$ and the vector (or the first dimension of tensor) $\bm{\mathcal{X}(x)}$, while $\bm{U}\bm{\mathcal{X}(x)}$ is a matrix-vector multiplication.
+The matrix $\bm{U}$ is a stack of nodal values $\bm{u}^{(j)}$, while $\bm{\mathcal{X}(x)}$ is a vectorized function of $\bm{x}$ that is parameterized with nodal coordinates $\bm{x}^{(j)}$ in the message passing operation.
+The graph node features (i.e., coordinate ($\bm{x}^{(j)}$) and value ($\bm{u}^{(j)}$)) are trainable parameters of the INN.
 If the nodal coordinates ($\bm{x}^{(j)}$) are fixed, one can find nodal values ($\bm{u}^{(j)}$) without changing the discretization of the input domain.
-If the nodal coordinates are also updated, the optimization will adjust the domain discretization similar to the r-adaptivity in FEM [23].
-Once the forward propagation is defined, the loss function is chosen based on the problem type: training, solving, and calibration.
+If the nodal coordinates are also updated, the optimization will adjust the domain discretization similar to r-adaptivity in FEM [21, 22].
+Once the forward propagation is defined, the loss function is chosen based on the problem type: training, solving, or calibrating.
 Special case 1: regular mesh with Tucker decomposition
 When the input domain is discretized with a regular mesh, we can significantly reduce the trainable parameters (or degrees of freedom, DoFs) by leveraging Tensor Decomposition (TD) [13].
-One of the widely used TD methods is the Tucker decomposition [37, 38].
+One of the widely used TD methods is Tucker decomposition [36, 37].
+It approximates a high-order tensor with $J_{i}$ nodes in $i$-th dimension as a tensor contraction between dimension-wise matrices and a core tensor $\bm{\mathcal{G}}$, which has the same order of the original tensor but with smaller nodes $M_{i}$ ($<J_{i}$).
+The $M_{i}$ is often called a "mode" to distinguish it from the original node $J_{i}$.
 Consider a three-input ($I=3$) and one output ($L=1$) system and assume the input domain is discretized with $J=J_{1}\times J_{2}\times J_{3}$ nodes, as shown in the special case box of Figure 1.
-To facilitate tensor notation, the nodal values $\bm{u}$ will be denoted with left/right super/sub scripts, $\prescript{m}{i}{u}^{(j)}_{l}$, where $i\in\mathbb{N}^{I}$ is the input index, $m\in\mathbb{N}^{M_{i}}$ is the mode index, $l\in\mathbb{N}^{L}$ is the output index and $j\in\mathbb{N}^{J_{i}}$ is the nodal index.
+To facilitate tensor notation, the nodal values will be denoted with left/right super/sub scripts, $\prescript{m}{i}{u}^{(j)}_{l}$, where $i\in\mathbb{N}^{I}$ is the input index, $m\in\mathbb{N}^{M_{i}}$ is the mode index, $l\in\mathbb{N}^{L}$ is the output index, and $j\in\mathbb{N}^{J_{i}}$ is the nodal index.
 The interpolated field $\mathcal{J}u(\bm{x})\in\mathbb{R}^{L=1}$ can be represented as a Tucker product:
 $\mathcal{J}u(\bm{x})=\llbracket\bm{\mathcal{G}};(\mathcal{J}\prescript{}{1}{\bm{u}})^{T},(\mathcal{J}\prescript{}{2}{\bm{u}})^{T},(\mathcal{J}\prescript{}{3}{\bm{u}})^{T}\rrbracket=\bm{\mathcal{G}}\times^{2}_{1}(\mathcal{J}\prescript{}{1}{\bm{u}})^{T}\times^{2}_{2}(\mathcal{J}\prescript{}{2}{\bm{u}})^{T}\times^{2}_{3}(\mathcal{J}\prescript{}{3}{\bm{u}})^{T},$ (4)
-where the core tensor $\bm{\mathcal{G}}\in\mathbb{R}^{M_{1}\times M_{2}\times M_{3}}$ is a full trainable matrix typically smaller than the original nodal discretization ($M_{1}<J_{1},M_{2}<J_{2},M_{3}<J_{3}$).
-This enables data compression.
-The $\mathcal{J}\prescript{}{i}{\bm{u}}(x_{i})\in\mathbb{R}^{M_{i}\times 1}$ is the 1-dimensional (1D) interpolated output of $i$-th input dimension over $M_{i}$ modes represented as:
-$\mathcal{J}\prescript{}{i}{\bm{u}}(x_{i})=\begin{bmatrix}\prescript{1}{i}{\bm{u}}^{(1)}&\prescript{1}{i}{\bm{u}}^{(2)}&\cdots&\prescript{1}{i}{\bm{u}}^{(J_{i})}\\ \prescript{2}{i}{\bm{u}}^{(1)}&\prescript{2}{i}{\bm{u}}^{(2)}&\cdots&\prescript{2}{i}{\bm{u}}^{(J_{i})}\\ \vdots&\vdots&\vdots&\vdots\\ \prescript{M_{i}}{i}{\bm{u}}^{(1)}&\prescript{M_{i}}{i}{\bm{u}}^{(2)}&\cdots&\prescript{M_{i}}{i}{\bm{u}}^{(J_{i})}\end{bmatrix}\cdot\begin{bmatrix}N^{(1)}(x_{i})\\ N^{(2)}(x_{i})\\ \vdots\\ N^{(J_{i})}(x_{i})\end{bmatrix}=\prescript{}{i}{\bm{U}}\times^{1}_{2}\prescript{}{i}{\bm{\mathcal{X}(x)}}=\prescript{}{i}{\bm{U}}\prescript{}{i}{\bm{\mathcal{X}(x)}},$ (5)
-where $\prescript{}{i}{\bm{U}}\in\mathbb{R}^{M_{i}\times J_{i}}$ and $\prescript{}{i}{\bm{\mathcal{X}(x)}}\in\mathbb{R}^{J_{i}\times 1}$.
-When there is more than one output $(L>1)$, the interpolated field becomes a vector of $L$ elements:
+where the core tensor $\bm{\mathcal{G}}\in\mathbb{R}^{M_{1}\times M_{2}\times M_{3}}$ is a trainable full matrix typically smaller than the original tensor that compresses data.
+The $\mathcal{J}\prescript{}{i}{\bm{u}}(x_{i})\in\mathbb{R}^{M_{i}\times 1}$ is one-dimensional (1D) interpolated output of $i$-th input dimension over $M_{i}$ modes represented as:
+$\mathcal{J}\prescript{}{i}{\bm{u}}(x_{i})=\begin{bmatrix}\prescript{1}{i}{u}^{(1)}&\prescript{1}{i}{u}^{(2)}&\cdots&\prescript{1}{i}{u}^{(J_{i})}\\ \prescript{2}{i}{u}^{(1)}&\prescript{2}{i}{u}^{(2)}&\cdots&\prescript{2}{i}{u}^{(J_{i})}\\ \vdots&\vdots&\vdots&\vdots\\ \prescript{M_{i}}{i}{u}^{(1)}&\prescript{M_{i}}{i}{u}^{(2)}&\cdots&\prescript{M_{i}}{i}{u}^{(J_{i})}\end{bmatrix}\cdot\begin{bmatrix}N^{(1)}(x_{i})\\ N^{(2)}(x_{i})\\ \vdots\\ N^{(J_{i})}(x_{i})\end{bmatrix}=\prescript{}{i}{\bm{U}}\times^{1}_{2}\prescript{}{i}{\bm{\mathcal{X}(x)}}=\prescript{}{i}{\bm{U}}\prescript{}{i}{\bm{\mathcal{X}(x)}},$ (5)
+where $\prescript{}{i}{\bm{U}}\in\mathbb{R}^{M_{i}\times J_{i}}$ and $\prescript{}{i}{\bm{\mathcal{X}(x)}}\in\mathbb{R}^{J_{i}}$.
+As illustrated in Figure 1’s special case box, the message passing (blue arrow) only happens in the axial directions, yielding 1D interpolation functions: $N^{(J_{i})}(x_{i})$.
+When there are more than one outputs $(L>1)$, the interpolated field becomes a vector of $L$ elements:
 $\mathcal{J}\bm{u}(\bm{x})$ $=\left[\mathcal{J}u_{1}(\bm{x}),\mathcal{J}u_{2}(\bm{x}),\cdots\mathcal{J}u_{L}(\bm{x})\right],\textit{where}$ (6)
-$\mathcal{J}u_{l}(\bm{x})$ $=\llbracket\bm{\mathcal{G}}_{l};(\mathcal{J}\prescript{}{1}{\bm{u}}_{l})^{T},(\mathcal{J}\prescript{}{2}{\bm{u}}_{l})^{T},(\mathcal{J}\prescript{}{3}{\bm{u}}_{l})^{T}\rrbracket=\bm{\mathcal{G}}_{l}\times^{2}_{1}(\mathcal{J}\prescript{}{1}{\bm{u}}_{l})^{T}\times^{2}_{2}(\mathcal{J}\prescript{}{2}{\bm{u}}_{l})^{T}\times^{2}_{3}(\mathcal{J}\prescript{}{3}{\bm{u}}_{l})^{T}.$
+$\mathcal{J}u_{l}(\bm{x})$ $=\llbracket\bm{\mathcal{G}}_{l};(\mathcal{J}\prescript{}{1}{\bm{u}}_{l})^{T},(\mathcal{J}\prescript{}{2}{\bm{u}}_{l})^{T},(\mathcal{J}\prescript{}{3}{\bm{u}}_{l})^{T}\rrbracket=\bm{\mathcal{G}}_{l}\times^{2}_{1}(\mathcal{J}\prescript{}{1}{\bm{u}}_{l})^{T}\times^{2}_{2}(\mathcal{J}\prescript{}{2}{\bm{u}}_{l})^{T}\times^{2}_{3}(\mathcal{J}\prescript{}{3}{\bm{u}}_{l})^{T},$
+where $\mathcal{J}\prescript{}{i}{\bm{u}}_{l}\in\mathbb{R}^{M_{i}\times 1}$.
+The trainable parameters are the core tensors $\mathcal{G}_{l}\in\mathbb{R}^{M_{1}\times M_{2}\times M_{3}},l=1,\dots,L$, and the nodal values $\prescript{}{i}{\bm{U}}_{l}\in\mathbb{R}^{M_{i}\times J_{i}},l=1,\dots,L$, yielding a total count of $L\left(\prod^{I}_{i}{M_{i}}+\sum^{I}_{i}{M_{i}J_{i}}\right)$ that scales linearly with the nodal discretization $J_{i}$.
 Special case 2: regular mesh with CP decomposition
 Tucker decomposition can be further simplified to CANDECOMP/PARAFAC (CP) decomposition [38, 39, 40] by setting the core tensor $\bm{\mathcal{G}}$ as an order-$I$ super diagonal tensor: $\bm{\mathcal{G}}\in\mathbb{R}^{M^{I}}$, $M=M_{1}=\cdots=M_{I}$, all zero entries except the diagonal elements.
-If we further set the diagonal elements of $\bm{\mathcal{G}}$ to be 1, the CP decomposition yields:
-$\mathcal{J}\bm{u}(\bm{x})$ $=\llbracket\bm{\mathcal{G}};(\mathcal{J}\prescript{}{1}{\bm{u}})^{T},(\mathcal{J}\prescript{}{2}{\bm{u}})^{T},(\mathcal{J}\prescript{}{3}{\bm{u}})^{T}\rrbracket=\sum^{M}_{m=1}{\prod^{I}_{i=1}{\mathcal{J}\prescript{m}{i}{u}(x_{i})}},$ (7)
-an inner product of the three interpolated vectors of length $M$.
-It is important to note that the interpolation function $N^{(j)}(x_{i})$ in Eq.5 is a one-dimensional function.
-In other words, both Tucker and CP decomposition replace a high-dimensional interpolation with multiple one-dimensional interpolations that dramatically reduce the number of trainable parameters (or DoFs).
+If we further set the diagonal elements of $\bm{\mathcal{G}}$ to be 1, the Tucker decomposition in Eq.
+6 becomes:
+$\mathcal{J}\bm{u}(\bm{x})=\sum^{M}_{m=1}{[\mathcal{J}\prescript{m}{1}{\bm{u}}(x_{1})\odot\dots\odot\mathcal{J}\prescript{m}{I}{\bm{u}}(x_{I})]},$ (7)
+where $\mathcal{J}\prescript{m}{i}{\bm{u}}(x_{i})\in\mathbb{R}^{L}$ and $\odot$ represents element-wise multiplication.
+In CP decomposition, the core tensor is no longer trainable, thus the total trainable parameter becomes $ML\sum^{I}_{i}{J_{i}}$.
+It is important to note that the interpolation function $N^{(j)}(x_{i})$ in Eq.5 is one-dimensional.
+In other words, both Tucker and CP decomposition replace a high-dimensional interpolation with one-dimensional interpolations that dramatically reduce the number of trainable parameters (or DoFs).
 As given in Table 1, the number of trainable parameters of the full interpolation scales exponentially with the discretization $J_{i}$, whereas the Tucker and the CP decomposition scale linearly with $J_{i}$.
-Table 1: Number of trainable parameters (or degrees of freedom, DoFs) of full interpolation, Tucker decomposition, and CP decomposition for an $I$-input single output relationship.
+Considering the trainable parameters of multi-layer perception (MLP) scales quadratically with the number of hidden neurons, INNs with tensor decomposition (TD) may dramatically reduce the model complexity and computing requirements.
+Table 1: Number of trainable parameters (or degrees of freedom, DoFs) of full interpolation, Tucker decomposition, and CP decomposition for an $I$-input single output $(L=1)$ relationship.
 The input domain is discretized with a regular grid of $J_{1}\times J_{2}\times\cdots\times J_{I}$ discretization.
-Assume that the nodal coordinates $(\bm{x}^{j})$ are fixed.
+We assume that the nodal coordinates $(\bm{x}^{(j)})$ are fixed.
 Full interpolation Tucker decomposition CP decomposition
@@
 The INN trainer optimizes nodal values $\bm{U}=\{\bm{u}^{(j)}\}_{j=1,\cdots,J}$ (and if needed, nodal coordinates ($\bm{X}=\{\bm{x}^{(j)}\}_{j=1,\cdots,J}$)) under a given loss function and training data.
-Consider a regression problem with $K$ labeled data: $(\bm{x}^{*}_{k},\bm{u}^{*}_{k})$, $k=1,\cdots,K$ and $\bm{x}^{*}_{k}\in\mathbb{R}^{I},\bm{u}^{*}_{k}\in\mathbb{R}^{L}$.
-Here the superscript * denotes data.
+Consider a regression problem with $K$ labeled data: $(\bm{x}^{*}_{k},\bm{u}^{*}_{k})$, $k=1,\cdots,K$; and $\bm{x}^{*}_{k}\in\mathbb{R}^{I},\bm{u}^{*}_{k}\in\mathbb{R}^{L}$.
+The superscript * denotes data.
 A mean squared error (MSE) loss function for this regression problem is defined as:
 $loss(\bm{U},\bm{X})=\frac{1}{K}\sum_{k}(\mathcal{J}\bm{u}(\bm{x}^{*}_{k})-\bm{u}^{*}_{k})^{2}.$ (8)
-Finally, an optimization problem is formulated as:
+Finally, an optimization is formulated as:
 $\underset{\bm{U},\bm{X}}{\text{minimize}}\ loss(\bm{U},\bm{X}).$ (9)
-Several benchmark problems are introduced in the Supplementary Information.
-The superior training accuracy, faster training and inference time, and fewer trainable parameters of the INN trainer are demonstrated.
+Section 3.1 exhibits the application of INN trainer to metal additive manufacturing while other benchmarks are elucidated in SI Section 2.
 INN solver
-INN solver generalizes classical numerical methods such as FEM and meshfree, as well as the model order reduction methods such as proper generalized decomposition (PGD) [42] and tensor decomposition (TD) [17, 43].
-Here, we introduce a formulation with CP decomposition to solve the generalized space $(\bm{x})$ time $(t)$ parameter $(\bm{\theta})$ PDE, which is prohibitive for most numerical methods and machine learning approaches [42].
+An INN solver generalizes classical numerical methods such as FEM and meshfree, as well as the model order reduction methods such as proper generalized decomposition (PGD) [41] and tensor decomposition (TD) [14, 42].
+Here, we introduce a formulation with CP decomposition to solve the generalized space $(\bm{x})$ time $(t)$ parameter $(\bm{\theta})$ (S-T-P) PDE, whose computational cost is prohibitively high for most numerical methods and machine learning approaches [41].
 Consider a parameterized space-time PDE:
-$\mathcal{L}u(\bm{x},t,\bm{\theta})=0,$ (10)
+$\mathcal{L}\bm{u}(\bm{x},t,\bm{\theta})=0,$ (10)
 where $\mathcal{L}$ is the general partial differential operator (can be linear or nonlinear).
-The INN solver uses the same neural network structure as the trainer, but the definition of loss function differs since training data is not known before solving the equation.
-Instead, the loss function is formulated as the weighted summation of the PDE residual [44, 45, 46].
-INN approximates the general space $(\bm{x})$ time $(t)$ parameter $(\bm{\theta})$ solution using the CP decomposition format:
-$\begin{split}\mathcal{J}{\bm{u}}(\bm{x},t,\bm{\theta};\bm{U},\bm{X})&=\sum_{m=1}^{M}[\mathcal{J}\prescript{m}{x_{1}}{\bm{u}}(x_{1})]\odot\cdots\odot[\mathcal{J}\prescript{m}{x_{d}}{\bm{u}}(x_{d})]\odot[\mathcal{J}\prescript{m}{t}{\bm{u}}(x_{t})]\\ &\odot[\mathcal{J}\prescript{m}{\theta_{1}}{\bm{u}}(\theta_{1})]\odot\cdots\odot[\mathcal{J}\prescript{m}{\theta_{k}}{\bm{u}}(\theta_{k})].\end{split}$ (11)
+The INN solver uses the same neural network structure as the trainer, but the loss function varies with the equations to be solved.
+It is formulated as the weighted summation of the PDE residual [43, 44, 45].
+An INN solution field in the S-T-P domain can be written as:
+$\begin{split}\mathcal{J}{\bm{u}}(\bm{x},t,\bm{\theta};\bm{U},\bm{X})&=\sum_{m=1}^{M}[\mathcal{J}\prescript{m}{x_{1}}{\bm{u}}(x_{1})]\odot\cdots\odot[\mathcal{J}\prescript{m}{x_{d}}{\bm{u}}(x_{d})]\odot[\mathcal{J}\prescript{m}{t}{\bm{u}}(x_{t})]\\ &\odot[\mathcal{J}\prescript{m}{\theta_{1}}{\bm{u}}(\theta_{1})]\odot\cdots\odot[\mathcal{J}\prescript{m}{\theta_{k}}{\bm{u}}(\theta_{k})],\end{split}$ (11)
 where $d$ and $k$ are the spatial dimension and the number of parameters, respectively.
-The operator $\odot$ refers to the element-wise multiplication.
 Similar to the trainer, the goal is to find nodal values $\bm{U}$ (and nodal coordinates $\bm{X}$, if one wants to adapt the mesh).
-As a result, INN obtains the space-time-parameter solution by minimizing the loss function:
+As a result, the INN obtains the S-T-P solution by minimizing the loss function:
 […]
 where $\delta\bm{u}$ is the weight function and can be defined using Galerkin, Petrov-Galerkin, collocation or quadratic formulation.
-It should be noted that INN should automatically satisfy Dirichlet and initial conditions.
-As a result, only weighted sum residual residual is considered in the loss function Eq.
-11.
-See the Supplementary Information for detailed derivations of the loss function.
-In Supplementary Information, a 3D linear elasticity problem is solved with the full interpolation INN and CP decomposition INN, and compared with FEM.
-Especially the INN with CP decomposition enables a voxel mesh of 10 million cube - $10^{21}$ grid points, which expands the state-of-the-art FEM limit by order of 10.
+Due to the Kronecker delta property of INN interpolation functions, Dirichlet boundary conditions and initial conditions can be strongly imposed.
+As a result, only the weighted sum residual is considered in Eq.
+11 as a loss function.
+This distinguishes INNs from most data-driven PDE solvers that weakly impose those conditions [46].
+See the SI Section 4 for detailed derivations of the loss function and solution scheme.
+INN solvers applied to the metal additive manufacturing and the linear elastic solid mechanics problem are introduced in Section 3.3 and SI Section 4, respectively.
 INN calibrator
@@
 where $\mathcal{J}\bm{u}(\bm{x})$ is the trained or solved forward model, $\bm{u}^{*}_{k}$ is the $k$-th measured observation and $\bm{x}^{*}$ is the calibrated input variable.
-A few calibration benchmarks are provided in Supplementary Information that demonstrate the superior calibration accuracy and reliability of the INN calibrator.
-3 Scientific and Engineering applications
-Figure 2: A summary of engineering applications of the INN methods to scientific and engineering problems.
-The calibration problem develops a reduced-order model of laser powder bed additive manufacturing to calibrate the heat source parameters from experimental data.
-A space-time-parameter solution of the same problem with large parameter space is demonstrated in Solving category.
-INN solvers result in significant storage reduction and fast simulation.
-Finally, a data-driven real-time online monitoring and feedback control tool is formulated with INN that uses only 18% training parameters and is 18-31 times faster compared to MLP.
-This section will illustrate how Engineering Software 2.0 enabled with INN can be applied to model various facets of metal additive manufacturing by integrating training, solving, and calibration processes.
-Note that the INN with CP decomposition will be referred to INN for brevity.
+The INN calibrator applied to the heat source calibration task in metal additive manufacturing can be found in Section 3.3 while other benchmarks can be found in SI Section 3.
+3 Scientific and Engineering applications: Metal Additive Manufacturing
+Figure 3: A summary of engineering applications of INNs to additive manufacturing (AM) problems.
+First, a data-driven real-time online monitoring and feedback control tool is formulated with an INN that uses only 18% training parameters of MLP and is 18-31 times faster than MLP.
+The calibration problem develops a reduced-order model of laser powder bed fusion (L-PBF) AM to calibrate the heat source parameters from experimental data.
+Finally, the INN solver solves a space-time-parameter heat transfer equation, resulting in a significant storage reduction and faster simulations.
+This section illustrates how Engineering Software 2.0 enabled with INNs can be applied to advance various facets of metal additive manufacturing (AM).
+Note that the INN with CP decomposition will be referred to as INN in this section for brevity.
 Laser powder bed fusion (L-PBF) is a type of metal additive manufacturing in which metal powders are layered and melted by a laser beam to create objects that match the provided CAD designs [47].
-The added flexibility of L-PBF comes with a very large design space and spatially varying microstructure after manufacturing.
-Therefore, considerable research is being devoted to employing in situ monitoring data for real-time control of L-PBF processes to minimize the variability [8].
-Computational modeling of L-PBF spans the simulation of the manufacturing process with a high dimensional design space (solving), computing the internal relationship among different manufacturing parameters (calibrating), and designing an online monitoring and control system for L-PBF (training).
-The governing equation for modeling the manufacturing process is given by
-$\frac{\partial T(\mathbf{x},t)}{\partial t}-k\Delta T(\mathbf{x},t)=Q_{h}(\mathbf{x},t,P,R,\eta,d)$ (14)
-This is a heat conduction equation where the moving laser heat source is incorporated through the source term, $Q_{h}$.
-The equation involves spatial parameter $\mathbf{x}$, temporal variable $t$, thermal conductivity $k$, power $P$, beam radius $R$, absorptivity $\eta$, and laser depth $d$.
-For the simplest case of single track laser pass, the initial condition is $T(\mathbf{x},0)=T_{0}(\mathbf{x})$, and the boundary conditions (BC) are $T=\tilde{T}$ in $\partial\Omega_{T}$ for Dirichlet BC and $\frac{\partial T}{\partial\mathbf{x}}\cdot\mathbf{n}=\bar{q}$ in $\partial\Omega_{q}$ for Neumann BC, where the spatial domain boundary $\partial\Omega=\partial\Omega_{T}\cup\partial\Omega_{q}$.
+The added flexibility of L-PBF comes with a huge design space and spatially varying microstructure after manufacturing.
+Therefore, considerable research is devoted to employing in situ monitoring data for real-time control of L-PBF processes to minimize the variability and optimize the target properties [48].
+Computational modeling of L-PBF spans the simulation of the manufacturing process with a high dimensional design space (solving), the identification of model parameters from sparse experimental data (calibrating), and the online monitoring and control of the L-PBF (training).
+The governing equation for modeling the manufacturing process is given by a heat conduction equation with a moving laser heat source $S_{h}$:
+$\frac{\partial T(\mathbf{x},t)}{\partial t}-k\Delta T(\mathbf{x},t)=S_{h}(\mathbf{x},t,P,\eta,d)$ (14)
+where the equation involves spatial variables $\mathbf{x}$, temporal variable $t$, and four variable parameters: thermal conductivity $k$, laser power $P$, absorptivity $\eta$, and laser depth $d$.
+For the simplest case of single track laser pass, the initial condition is $T(\mathbf{x},0)=T_{0}(\mathbf{x})$, and the boundary conditions (BCs) are $T=\tilde{T}$ in $\Gamma^{D}$ for Dirichlet BC and $\frac{\partial T}{\partial\mathbf{x}}\cdot\mathbf{n}=\bar{q}$ in $\Gamma^{N}$ for Neumann BC, where the spatial domain boundary $\Gamma=\Gamma^{D}\cup\Gamma^{N}$.
+3.1 Real-time Online Control of Laser Powder Bed Fusion Additive Manufacturing
+The first application is fully data-driven, where the data are generated from numerical methods and the INN trainer is used to develop an online control system for L-PBF AM.
+The goal is to maintain a homogeneous temperature of the melt pool (i.e., a molten region near the laser spot) across each layer using real-time sensor data, under the hypothesis that a homogeneous melt pool temperature across each layer of material deposition will reduce the variability in the microstructure and mechanical properties.
+The laser power at each point of a certain layer needs to be controlled (increased/decreased by certain amount compared to previous layer) to achieve this homogeneity in melt pool temperature.
+The fundamental challenge of implementing a model predictive control system for such an application is the computational resources required for forward prediction (involving finite element or computational fluid dynamics models) and inverse prediction, as illustrated in Figure 3.
+The INN trainer aims to provide a reliable and memory-efficient surrogate model that can replace mechanistic models in predictive control systems.
+In this application, the database is generated by fusing experiments and computational methods (see SI Section 2.6 for details) for aluminum alloy.
+The online control loop uses the trained INN models to inform the manufacturing machine.
+There are two models: forward model and inverse model.
+The forward model uses the mechanistic features and the thermal emission plank (TEP, a representative of the melt pool temperature) to predict the TEP of the next layer of the build when no correction is applied.
+The mechanistic features are precomputed based on the tool path and geometry of the build to generalize a data-driven model for unseen tool path and geometric features [49]).
+The target TEP for the next layer is the average of the forward model’s outputs (i.e., predicted TEPs for the next layer) measured at the current layer.
+The inverse model takes this target TEP as the input along with the previous layer’s mechanistic features to predict the required laser power for the next layer.
+Finally, the control software calculates how much change in laser power is required at each point, and sends the information to the machine.
+The corrected laser power is then used to print the next layer.
+The forward model is trained on experimentally observed data, while the inverse model is trained on computational data coming from a finite-difference solver [18].
+The INN model is applied for both forward and inverse prediction and is compared with the MLP.
+Figure 3 shows that for the inverse model INN is at least 18 to 31 times faster to train compared to the MLP to reach the same level of training error, and it reduced the number of trainable parameters by a staggering amount of 82%.
+The speed-up depends on the amount of data used.
+The INN method is consistently more accurate compared to MLP (see, SI Section 2.6 for more details).
 3.2 Calibration of Heat Source Parameters for Additive Manufacturing
-In this example, the INN is used for calibrating a model of heat source of L-PBF, which uses a laser heat source to selectively melt metal powder bed to form a connected shape.
-A commonly used Gaussian beam profile can be modeled as a volumetric heat source [48].
-The heat source equation reads,
-$Q_{h}(\textbf{x},t,\textbf{p})=\frac{2\mathcal{P}\eta}{\pi r^{2}_{b}d}e^{-\frac{2(x^{2}+y^{2})}{r^{2}_{b}}}$ (15)
-where, $z_{top}-z\leq d$, $z_{top}$ is z coordinate of the current powder bed surface.
-The notations follow Eq.
-14 and $r_{b}$ is the standard deviation of the Gaussian source.
-To get an accurate representation of the physics behind it, model parameters need to be calibrated against experimental data.
-Following [49], three calibration parameters $p_{1}$, $p_{2}$, and $p_{3}$ can be defined as, $d=p_{1}\frac{\mathcal{P}}{V}RHF^{2}$, $\eta=p_{2}\frac{\mathcal{P}}{V}RHF^{2}$, and $r_{b}=p_{3}\frac{\mathcal{P}}{V}RHF^{2}$.
-Here, $P$ is laser power, $V$ is the scan speed, and $RHF$ is a residual heat measure predetermined based on a laser toolpath.
-The challenges for calibrating these parameters are: 1) the computational models are often very expensive to perform even one forward run, 2) the computational models are not automatically differentiable, and 3) limited experimental data is available.
-As a demonstration for this problem, INN with 1760 trainable parameters is trained on 44 sets of meltpool temperature data generated by the finite difference method (FDM) using 4 different power-scanning speed setup.
-Later, this trained INN is used for optimizing the three aforementioned parameters using gradient-based optimizers.
-To show the superiority of INN over traditional neural networks in such data-scarce applications, a 3-layer MLP with 2689 trainable parameters is trained using the same data fed to INN.
-Details regarding this calibration process are provided in the Supplementary Information.
-Both INN-based and MLP-based calibrations are repeated 50 times to give a statistical comparison (see Figure 2).
-Even with only 44 sets of simulation data, the INN calibrated parameters that produce mean meltpool temperature (or TEP, thermal emission plank) within 7.9% of the experimental data.
-On the contrary, MLP produced a difference of 44.1% against experimental data, which shows that it hardly learns the underlying physics using such small data with even more trainable parameters.
+In this example, an INN calibrates the heat source function $S_{h}$ in Eq.14 for experimental data. A Gaussian beam profile is modeled as a volumetric heat source [50], which is written as:
+$S_{h}(\mathbf{x},t,\mathbf{p})=\frac{2P\eta}{\pi r^{2}_{b}d}e^{-\frac{2(x^{2}+y^{2})}{r^{2}_{b}}}\cdot\mathbf{1}_{\{z|z_{top}-z\leq d\}}(z)$ (15)
+where $z_{top}$ is the z-coordinate of the current surface of the powder bed and $\mathbf{1}$ is the indicator function.
+The notation follows Eq.
+14, and $r_{b}$ is the Gaussian profile’s standard deviation that controls the beam’s width.
+Following reference [10], the heat source parameters $d,\eta,r_{b}$ are controlled by the calibration parameters $p_{1}$, $p_{2}$, $p_{3}$ such that $d=p_{1}\frac{P}{V}RHF^{2}$, $\eta=p_{2}\frac{P}{V}RHF^{2}$, and $r_{b}=p_{3}\frac{P}{V}RHF^{2}$.
+Here, $V$ is the laser scan speed and $RHF$ is a predetermined residual heat factor based on the laser toolpath [51].
+This calibration problem’s major challenge is achieving a highly accurate surrogate model given a sparse set of training data.
+INNs outperform MLPs in solving these data-scarce problems because they are rooted in the interpolation theory.
+As a demonstration, an INN with 1760 trainable parameters is trained on 28 sets of meltpool temperature data generated by the finite difference method (FDM) [18] using 4 different power-scanning speed set-ups.
+Later, this trained INN is used as a surrogate to optimize the three parameters ($p_{1},p_{2},p_{3}$) given the experimental data.
+To show the superiority of INNs over traditional neural networks in such data-scarce applications, a 3-layer MLP with 2689 trainable parameters (roughly the same order of magnitude as the INN) is trained using the same data fed to the INN.
+In practice, we observed that INN converges well within 100 epoch, while MLP requires over 1000 epoch to converge to a test MSE 10 times larger than INN given all other setup the same.
+Details on this training and calibration process are provided in SI Section 3.1.
+The INN- and MLP-based calibrations are repeated 50 times to provide a statistical comparison (see Figure 3).
+Even with only 28 sets of simulation data, the calibrated parameters using the INN produce a mean melt pool temperature within a 7.9% difference of the experimental data.
+In contrast, the parameters calibrated from the MLP surrogate produce a difference of 44.
+1%, which is unsatisfactory.
+This is expected since MLP is a general model requiring a large amount of data to achieve a reliable representation of the problem.
+On the other hand, the INN achieves a much lower difference, showing that the model can learn underlying physics even with a scarce dataset.
 3.3 Solving High Dimensional Space-Time-Parameter Heat Equation
-The governing equation Eq.14 has 8 input variables; 3 for spatial domain ($x,y,z$), 1 for time domain ($t$), and 4 for parametric domain ($k,p,e,d$).
-See Supplementary Information for a detailed problem definition.
-This equation is not feasible for process design with most existing computational methods, including the data-driven method inspired by the finite element method (FEM) and the physics-informed neural network (PINN) [20].
-To solve the problem with FEM-based data-driven method, one first needs to sample from the 4-dimensional parametric domain, followed by running the time-stepping-based finite element analysis for each sampled parameter set.
-Then the FEM solutions should be stored as offline training data, and a neural network surrogate model is trained using the data.
-The computational cost for such operations is prohibitive since large space-time mesh is required in each simulation for real engineering application, as detailed in the Supplementary Information.
-Moreover, as a larger number of training data is generated due to the high dimensionality of the input domain, this method becomes impractical in terms of data storage.
-Furthermore, running a large-scale FEM simulation requires a large amount of computational memory.
-PINN is a unified framework of training and solving for such parameterized PDEs; however, the computational cost still grows exponentially as the number of input variables grows.
-As illustrated in Supplementary Information, it is very challenging to solve this 8-dimensional equation.
-In addition, the convergence of solution accuracy is not guaranteed especially for such high-dimensional problems.
-Thus, it remains to be seen if PINN can even find a good solution for more than 7 input variables.
-PINN also suffers from memory issues for GPU training due to the exponential growth of the collocation points although batch training can be adopted to alleviate memory usage.
-Contrary to FEM and PINN, INN solves this space-time-parameter problem all-at-once.
-As it is based on the projected 1D interpolation, the number of trainable parameters increases linearly with the input variables, making the computational cost manageable.
-Meanwhile, since tensor-decomposed data structures are leveraged, INN has the benefits of significant memory efficiency and storage gain, especially for high-dimensional problems.
-In Figure 2, we solve the 8-dimensional space-time-parameter PDE with the total number of space-time-parameter grid points of $3\times 10^{32}$, where we use 25,000 grid points to discretize each spatial and parametric coordinate and 50 grid points for the temporal coordinate.
-The details of the solver example are provided in the Supplementary Information.
-Since the number of trainable parameters of INN grows linearly, INN has a significant storage gain of $1.7\times 10^{25}$ compared to the classical FEM.
-Moreover, when using INN, once the parametric solution of the AM problem is obtained, one can easily investigate the uncertainty propagation by imposing probabilistic distributions on the inputs of INN, thanks to the tensor-decomposed data structure.
-Details on the statistical prediction of parametric solution can be found in the Supplementary Information.
-3.3 Real-time Online Control of Laser Powder Bed Fusion Additive Manufacturing
-In this application, the INN trainer is used to devise an online control system for L-PBF additive manufacturing.
-The goal is to maintain a homogeneous melt pool temperature across each layer.
-The hypothesis is that a homogeneous melt pool temperature across each layer of material deposition will reduce the variability in the microstructure and mechanical properties.
-The fundamental challenge of implementing a model predictive control system is the computational time required for the forward prediction (involving finite element or computational fluid dynamics models).
-For prediction, the data is generated by fusing experiments and computational methods (see Supplementary Information for details).
-The online control loop uses the trained INN model to inform the manufacturing machine.
-There are two models: forward model and inverse model.
-The forward model uses the mechanistic features (precomputed based on the tool path and geometry of the build to generalize a data-driven model for unseen tool path and geometric features; more details can be found in [50]), and melt pool temperature of the previous layer (measured by the sensor as Thermal Emission Plank (TEP) signal; more details in the Supplementary Information) to predict the TEP/melt pool temperature in the next layer for an uncontrolled process.
-The uncontrolled TEP is converted to a homogenized TEP over the entire layer.
-The inverse model takes this targeted, homogenized TEP as the input along with the previous layer’s mechanistic features to predict the required laser power for the next layer.
-The laser power is then exported to the machine to print the next layer.
-The forward model is trained on experimentally observed data while the inverse model is trained on computational data coming from a finite element solver.
-The INN model is applied for both forward and inverse prediction and compared against MLP.
-Figure 2 shows that for the inverse model INN is at least 18 to 31 times faster to train compared to the MLP to reach the same level of training error, and it uses only about 18% of the parameters that MLP uses.
-The speed-up depends on the amount of data used.
-The INN method is consistently more accurate compared to MLP (see Supplementary Information for more details).
+[…]
+The surrogate model $f$ can also be obtained using the classical data-driven approach.
+In the offline stage, training data can be generated by repeatedly running numerical solvers (such as finite difference method (FDM) and variational multiscale method (VMS)) in the Space-Time domain with different sets of parameters $\mathbf{{p}}$.
+Then a surrogate model is trained with the data.
+The standard data-driven approach suffers from the curse of dimensionality when the parametric space is high-dimensional.
+As a result, the data-driven approaches face three challenges at the offline stage: expensive computation for repetitive data generation, huge memory cost for running full-scale simulations, and excessive disk storage of offline training data.
+We estimate the time and storage required for the data-driven approach with FDM and VMS, as shown in Table 2 and illustrated in Figure 2.
+Unlike the data-driven approaches, INN solvers treat the parameters $\mathbf{p}$ as additional parametric inputs.
+As a result, INNs can directly obtain a parametric surrogate model $f=u(\mathbf{x},t,\mathbf{p})$ directly from the governing equation without going through the cumbersome offline data generation and surrogate model training process.
+This is particularly beneficial to high-dimensional parametric surrogate models.
+Moreover, our results prove the INN solution (or the S-T-P surrogate model) achieves an exceptionally high accuracy of $R^{2}$ = 0.9969, considering that most data-driven approaches for S-T-P problems trained on numerical simulation data suffer from low model accuracy, typically below $R^{2}=0.9$ [52].
+It is worth mentioning that the physics-informed neural networks (PINNs) can handle the S-T-P problem similar to the INN solver.
+However, the number of collocation points scales exponentially with the number of inputs, and our preliminary study revealed that PINN fails to converge for the same 8-dimensional S-T-P problem after consuming considerable computational resources (see SI Section 4.4).
+The performance comparison of INN versus data-driven methods is shown in Table.
+2.
+Table 2: Performance comparison of different surrogate models.
+The INN’s discretization in the 4D parametric space is $100^{4}$, so the corresponding sampling points required for the data-driven methods should be $10^{8}$.
+The total simulation time and storage requirement of data-driven approaches are estimated from a single run.
+Detailed explanations for this benchmark can be found in SI Sections 4.3 and 4.4.
+Method
+Resolution
+($\mu$m)
+Number of
+offline simulations
+Total offline
+simulation time
+Storage
+Data-driven (FDM) 25 $1\times 10^{8}$ 2,600 years $7.82\times 10^{8}$
+Data-driven (VMS) 12.5 $1\times 10^{8}$ 9 years $8.79\times 10^{6}$
+INN 8 $1$ 15 minutes $7.37\times 10^{-2}$
 4 Conclusions
-We show that interpolating neural networks (INN) can train, solve, and calibrate scientific and engineering problems that are extremely challenging or prohibitive for existing numerical methods or machine learning models.
-The keys to INN’s success in those problems are 1) it interpolates virtual points with well-established interpolation theories (not interpolating training data with a black-box network) and 2) it leverages the idea of separation of variables and tensor decomposition to resolve the curse of dimensionality.
+This article demonstrates that the interpolating neural network (INN) can train, solve, and calibrate scientific and engineering problems that are extremely challenging or prohibitive for existing numerical methods and machine learning models.
+The keys to INN’s success are 1) it interpolates the graph nodes’ values with well-established interpolation theories (not interpolating training data with a black-box network that often causes overfitting issues) and 2) it leverages the tensor decomposition (TD) to resolve curse of dimensionality.
 Due to the significantly reduced trainable parameters without losing approximation capabilities, INN can train/solve/calibrate faster with less memory requirement, and achieve extremely high model accuracy and fast inference time.
-As this is the original article on INN, we have myriads of research questions that may gather huge research interests across various areas including but not limited to scientific machine learning, applied mathematics, computer vision, and data science.
+As this is the original article on INNs, we have myriads of research questions that may gather huge research interests across various areas including but not limited to scientific machine learning, applied mathematics, computer vision, and data science.
 To mention a few:
-  • Multi-resolution aspect (i.e., a varying number of segments across INN networks, $M$) of INN needs to be investigated. See Supplementary Information for further discussions.
-  • Convergence of INN for training can be proved mathematically.
-  • INN training on ensembles of training data might be a way to enhance the generality of the model for image problems or noisy data.
-  • Since INN could be used as both a solver for physics-based problems and a function approximator for data-driven problems, it is of interest to combine the two behaviors together into one single model, so that we can solve large scale problems involving both complex physics and scarce data.
-For example, the calibration task of LPBF heat source.
-We expect INN can open a next-generation Engineering Software 2.0 by theoretically integrating training, solving, and calibration and tackling extremely challenging problems in the science and engineering domain.
+  • Depending on the problem, INN’s superior training efficiency may cause overfitting issues.
+The ensemble training (i.e., training on random subsets of the training data) might be a way to enhance the generality of the model for image problems or noisy data.
+  • Multi-resolution aspect (i.e., a varying number of segments across TD modes) of INNs needs to be investigated. See SI 1.4 for further discussions.
+  • Similar to the INN solvers, the convergence of INN trainers can be studied mathematically.
+  • Since INNs can be used as both a solver for physics-based problems and a function approximator for data-driven problems, it is of interest to combine the two behaviors into one single model to solve large-scale problems involving both complex physics and scarce data.
+  • The superior performance of the INN solver may facilitate multiscale analysis in a vast parametric space.
+One can integrate the S-T-P INN solver with a concurrent multiscale analysis framework such as self-consistent clustering analysis [53, 54] and open a new direction towards parameterized multiscale analysis.
+  • The current INN code has been optimized to run on a single GPU. Since the INN forward pass over multiple modes can be parallelized, multi-GPU programming will further speed up the code.
+  • Although Tucker decomposition has been introduced as a special case 1 in Section 2, the numerical experiments conducted in this article mostly focus on the CP decomposition.
+Theoretically, Tucker decomposition has a larger approximation space than CP decomposition.
+Further studies on the numerical aspects of the Tucker decomposition are needed.
+We expect INNs will open a next-generation Engineering Software 2.0 that revolutionizes almost all fields of computational science and engineering.
+Acknowledgments
+We express our sincere gratitude to Dr.
+Gino Domel, Mr.
+Joseph P.
+Leonor, Mr.
+Stefan Knapik, and Mr.
+Vispi Karkaria from Northwestern University for their invaluable comments and support throughout this work.
+We also acknowledge the following grants: DEVCOM Army Research Laboratory under Cooperative Agreement Numbers W911NF-20-2-0292 and W911NF-21-2-02199 to J.
+Cao and the startup funds provided by Virginia Polytechnic Institute and State University to S.
+Saha.
+C.
+Park acknowledges the internship opportunity with the Freudenberg Group, within the team managed by Dr.
+Vishwanath Hegadekatte.

### v5

@@
-Interpolating neural network: A lightweight yet precise architecture for data training, equation solving, and parameter calibration
+Interpolating neural network: A novel unification of machine learning and interpolation theory
 Abstract
@@
 However, applying this transition in engineering software presents challenges, including low surrogate model accuracy, the curse of dimensionality in inverse design, and rising complexity in physical simulations.
-We introduce an interpolating neural network (INN), grounded in interpolation theory and tensor decomposition, to realize Engineering Software 2.0 by advancing data training, partial differential equation solving, and parameter calibration.
+We introduce an interpolating neural network (INN), grounded in numerical methods, to realize Engineering Software 2.0 by advancing data training, partial differential equation solving, and parameter calibration.
 INN offers orders of magnitude fewer trainable/solvable parameters for comparable model accuracy than traditional multi-layer perceptron (MLP) or physics-informed neural networks (PINN).
@@
 However, the application of these technologies in the fields of engineering and science presents unique challenges.
-One such challenge is the relatively low generalization accuracy of machine learning (ML)-based solvers compared to traditional numerical methods [4, 5, 6, 7, 8].
+One such challenge is the relatively low generalization accuracy of machine learning (ML)-based partial differential equation (PDE) solvers compared to traditional numerical methods [4, 5, 6, 7, 8].
 After in-depth research was conducted on those methods, an increasing number of researchers began to question their effectiveness [8, 9].
@@
 Furthermore, the complexity of physical simulations continues to increase.
-For instance, additive manufacturing (AM) or integrated circuit (IC) simulations necessitate extremely fine resolution with high accuracy and multi-physics/multi-scale analysis within a reasonable simulation time [10, 11, 12].
-In this article, we propose the concept of Engineering Software 2.0, which aims to advance the current generation of computational and data-driven engineering software into a new paradigm.
-“Engineering Software 2.0 is an end-to-end software system that unifies training, solving, and calibrating in science/engineering problems."
+For instance, additive manufacturing (AM) or integrated circuit (IC) simulations necessitate extremely fine resolution with high accuracy and multi-physics/multi-scale analysis [10, 11, 12].
+In this article, we propose the concept of Engineering Software 2.0, which aims to advance the current generation of numerical and data-driven engineering software into a new paradigm.
+“Engineering Software 2.0 is an end-to-end software system that unifies and advances data training, PDE solving, and parameter calibration in science/engineering problems."
 We introduce interpolating neural networks (INNs) to achieve the new paradigm of Engineering Software 2.0.
-The key ideas behind INNs are: 1) discretize an input domain into non-overlapping segments whose bounds are denoted by interpolation nodes, 2) construct a graph with the interpolation nodes and formulate the message passing operation as a form of interpolation functions, and 3) optimize the values and coordinates of the interpolation nodes for a given loss function.
-Figure 1 shows a pictorial description of the approach.
-When the input domain is discretized with a regular mesh (see the special case in Figure 1), an INN leverages tensor decomposition (TD) [13, 14, 15, 16] to convert the growth of the computational cost of a high-dimensional problem from exponential to linear.
-INN facilitates three major tasks in computational science and engineering: 1) trains on both experimental and simulation data with very high accuracy and minimal memory footprint, 2) solves high-dimensional parameterized partial differential equations (PDEs), and 3) calibrates uncertain parameters rapidly and accurately.
+INNs generalize numerical methods such as the finite element method (FEM) [13, 14, 15] as a subset of deep neural network (DNN) or graph neural network (GNN).
+Although different groups of researchers have developed the two (FEM and neural networks), they share one thing in common; the goal is to find a function.
+The former finds a solution field of a differential equation while the latter finds an input-output relationship.
+Although it is possible to formulate INN from a DNN, we adopt GNN architecture to explain INN as a special case of GNN with a set of rules (i.e., interpolation).
+GNNs provide a general formulation to process information available in a graph, i.e.
+a cloud of interconnected points or nodes [16, 17, 18].
+Each node receives information from its connected neighbors that is then processed by a feedforward neural network to make a new prediction at that node.
+However, their network architectures are often overly complex, making convergence difficult for highly nonlinear problems.
+In Science and Engineering, most computer simulations also occur in interconnected point clouds (or nodes), but each interconnected node is associated with simpler interpolatory functions that are fast to differentiate and easy to integrate, leading to methods that obey established error bounds and convergence rates.
+This provides confidence in the predictions made but also imposes a need for relatively fine discretizations (dense point clouds) because each point is related to its connected neighbors by a simple function (e.g., low order polynomials or splines).
+INNs bridge the gap between the extreme generality of GNNs and the stringent restrictions imposed by numerical methods by constructing interpolatory GNNs that can simplify to known numerical methods or complexify to known GNN architectures.
+INNs consist of three processes: 1) discretize an input domain into non-overlapping segments whose bounds are denoted by interpolation nodes, 2) construct a graph with the interpolation nodes and formulate the message passing operation as a form of interpolation functions, and 3) optimize the values and coordinates of the interpolation nodes for a given loss function.
+Figure 1 describes the approach.
+When the input domain is discretized with a regular mesh (see the special case in Figure 1), INNs can leverage tensor decomposition (TD) [19, 20, 21, 22] to convert the growth of the computational cost of a high-dimensional problem from exponential to linear.
+INNs facilitate three major tasks in computational science and engineering: 1) data training, 2) PDE solving, and 3) parameter calibration, by reducing trainable parameters, training time, and memory/storage requirements without compromising model accuracy.
 […]
@@
 A trained/solved INN model can then be employed as a forward model of a calibrator to solve an inverse problem.
-Figure 2: Benchmark comparison of the single-scale finite difference method (FDM) solver [18], the variational multiscale FEM solver [11], and the INN solver.
-The data points for the first two methods with dashed marker edges are estimated, while that of INN with a solid marker edge is computed.
-A 3D-space, time, and 4D-parameter heat transfer equation is solved to model a 10 mm single-track laser powder bed fusion (L-PBF) metal AM.
-Detailed problem definition and explanation can be found in Section 3.
-All solvers are graphics processing units (GPU) optimized with the JAX library [19].
-2 INN: Unification of Training/Learning, Solving, and Calibrating
-Interpolating Neural Network (INN) generalizes numerical methods such as the finite element method (FEM) [20, 21, 22] as a subset of deep neural network (DNN) via graph network operations.
-Although different groups of researchers have developed the two (FEM and DNN), they share one thing in common; the goal is to find a function.
-The former finds a solution field of a differential equation while the latter finds an input-output relationship.
-INN is built upon this similarity and forms the backbone of Engineering Software 2.0 that unifies training, solving, and calibrating, as illustrated in Figure 1.
+Figure 2: Illustration of a 1D input domain discretized with 5 segments and 6 nodes.
+2 Results
 Consider a regression problem that relates $I$ inputs and $L$ outputs.
@@
 Each segment occupies a subspace of the $I$-dimensional input domain.
-For instance in Figure 1, the Euclidean mesh consists of segments of hexahedrons and triangular pillars.
-Each node (e.g., node $j$) has two features: nodal coordinates $(\bm{x}^{(j)}\in\mathbb{R}^{I})$ and values $(\bm{u}^{(j)}\in\mathbb{R}^{L})$.
+Each node (e.g., node $j$) has features: nodal coordinates $(\bm{x}^{(j)}\in\mathbb{R}^{I})$ and values $(\bm{u}^{(j)}\in\mathbb{R}^{L})$.
 Note that the superscript with parentheses refers to the graph node index.
@@
 A typical message passing in a graph neural network (GNN) returns a hidden state for each graph node which is mostly a tensor (including matrix and vector) [16].
-In contrast, the INN message passing returns a function (i.e., interpolation function $N^{(j)}(\bm{x})$, or often called shape function in computational mechanics) for each node $(j=1,\dots,J)$ as a hidden state.
+In contrast, the INN message passing returns a function (i.e., interpolation function $N^{(j)}(\bm{x})$, or often called shape function in FEM) for each node $(j=1,\dots,J)$ as a hidden state.
 A general $Q$-step message passing of an INN, $\mathcal{M}^{[q]}(*)$, can be expressed as:
 $N^{[q](j)}(\bm{x})=\mathcal{M}^{[q]}\left(\bm{x}^{(j)},\{\bm{x}^{(k)},\bm{e}^{(j,k)}|k\in\mathcal{N}^{s}(j)\},N^{[q-1](j)}(\bm{x})\right),\quad q=1,\cdots,Q,\quad N^{[0](j)}(\bm{x})=0,$ (1)
-where $N^{[q](j)}(\bm{x})$ is the interpolation function at node $j$ after $q$-th message passing and $\mathcal{N}^{s}(j)$ is a set of neighboring nodes of the center node $j$ with $s$ connections (i.e., $s$-hops, see Supplementary Information (SI) Section 1.1 for visual illustrations).
+where $N^{[q](j)}(\bm{x})$ is the interpolation function at node $j$ after $q$-th message passing and $\mathcal{N}^{s}(j)$ is a set of neighboring nodes of the center node $j$ with $s$ connections (i.e., $s$-hops, see Figure 2 and Supplementary Information (SI) Section 1.1 for visual illustrations).
 It is worth noting that the interpolation functions satisfy the Kronecker delta property, i.e., $N^{[q](j)}(\bm{x}^{(k)})=\delta_{jk}$.
 The operation $\mathcal{M}^{[q]}(*)$ constructs an interpolation function $N^{[q](j)}(\bm{x})$ for a graph node $\bm{x}^{(j)}$ using neighboring nodal coordinates and edge information: ${\bm{x}^{(k)},\bm{e}^{(j,k)}}$, and the interpolation function of the previous message passing: $N^{[q-1](j)}(\bm{x})$.
-This can be any interpolation method such as Lagrange polynomials [28, 20], meshfree [29, 30, 31], and splines [32, 33].
-The $Q$-step message passing implies that it is also possible to progressively construct nonlinear or more problem-specific interpolation functions, theoretically based on the generalized finite element method [34, 35] and convolution hierarchical deep learning neural networks [7, 15].
+When there is only one message passing ($Q=1$) with $s=1$ hop, INNs degenerate to the standard FEM linear shape function (or piecewise linear interpolation).
+As illustrated in Figure 2 (left), FEM with linear elements is the most localized approximation of an INN.
+On the other hand, one can progressively enlarge the support domain and the approximation capability of an interpolation function by adding message passings with a higher hop.
+For instance, the second message passing with $s=2$ hop constructs compact-supported interpolation functions with higher nonlinearity.
+This is theoretically equivalent to the generalized finite element method (GFEM) [29] and convolution hierarchical deep learning neural networks (C-HiDeNN) [7, 21].
+The message passing with a max-hop (Figure 2, right) can even make it a non-local approximation where support of the interpolation functions occupy the global domain, mimicking meshfree methods that exhibit superconvergence (i.e., faster convergence rate than the complete order of polynomial basis) [30, 31, 32].
 Depending on the choice of interpolation technique, other hyper-parameters can be involved in $\mathcal{M}^{[q]}(*)$ such as the activation (or basis) function and dilation parameter [7, 21].
 See SI Section 1 for various choices of the message passing operation.
+Regardless of the choice of interpolation technique for the message passing, the global degrees of freedom (i.e., number of nodes) remain constant.
+This distinguishes INNs from higher-order FEM where each segment is enriched with additional nodes.
+INNs, on the other hand, adapt nodal connectivity through the adjacency matrix and basis functions to reproduce almost any interpolation technique available in numerical methods.
+The compact-supported interpolation functions enable INNs to be optimized locally, making INNs distinguishable from MLPs.
+Since an MLP is a global approximator, INNs with compact-supported interpolation functions 1) converge faster than MLPs with the same number of trainable parameters (see Section 2.1) and 2) facilitate training on a sparse dataset (see Section 3.2), although it might end up with overfitting.
 During the forward propagation, an input variable $\bm{x}\in\mathbb{R}^{I}$ enters each graph node’s interpolation function $N^{[Q](j)}(\bm{x})$, followed by a graph-level readout operation:
@@
 The readout operation can be written as a tensor contraction (or matrix multiplication):
-$\mathcal{J}\bm{u}(\bm{x})=\begin{bmatrix}|&|&|&|\\ \bm{u}^{(1)}&\bm{u}^{(2)}&\cdots&\bm{u}^{(J)}\\ |&|&|&|\end{bmatrix}\cdot\begin{bmatrix}N^{(1)}(\bm{x})\\ N^{(2)}(\bm{x})\\ \vdots\\ N^{(J)}(\bm{x})\end{bmatrix}=\bm{U}\times^{1}_{2}\bm{\mathcal{X}(x)}=\bm{U}\bm{\mathcal{X}(x)},$ (3)
+$\mathcal{J}\bm{u}(\bm{x})=\begin{bmatrix}|&|&|&|\\ \bm{u}^{(1)}&\bm{u}^{(2)}&\cdots&\bm{u}^{(J)}\\ |&|&|&|\end{bmatrix}\cdot\begin{bmatrix}N^{(1)}(\bm{x})\\ N^{(2)}(\bm{x})\\ \vdots\\ N^{(J)}(\bm{x})\end{bmatrix}=\bm{U}\bm{\mathcal{X}(x)},$ (3)
 where $\bm{u}^{(j)}\in\mathbb{R}^{L},\bm{U}\in\mathbb{R}^{L\times J},\bm{\mathcal{X}(x)}\in\mathbb{R}^{J}$.
-Here, $\bm{U}\times^{1}_{2}\bm{\mathcal{X}(x)}$ denotes the tensor contraction operation between the second dimension of tensor $\bm{U}$ and the vector (or the first dimension of tensor) $\bm{\mathcal{X}(x)}$, while $\bm{U}\bm{\mathcal{X}(x)}$ is a matrix-vector multiplication.
 The matrix $\bm{U}$ is a stack of nodal values $\bm{u}^{(j)}$, while $\bm{\mathcal{X}(x)}$ is a vectorized function of $\bm{x}$ that is parameterized with nodal coordinates $\bm{x}^{(j)}$ in the message passing operation.
@@
 If the nodal coordinates are also updated, the optimization will adjust the domain discretization similar to r-adaptivity in FEM [14, 15].
-Once the forward propagation is defined, the loss function is chosen based on the problem type: training, solving, or calibrating.
-Special case 1: regular mesh with Tucker decomposition
-When the input domain is discretized with a regular mesh, we can significantly reduce the trainable parameters (or degrees of freedom, DoFs) by leveraging Tensor Decomposition (TD) [13].
-One of the widely used TD methods is Tucker decomposition [36, 37].
-It approximates a high-order tensor with $J_{i}$ nodes in $i$-th dimension as a tensor contraction between dimension-wise matrices and a core tensor $\bm{\mathcal{G}}$, which has the same order of the original tensor but with smaller nodes $M_{i}$ ($<J_{i}$).
-The $M_{i}$ is often called a "mode" to distinguish it from the original node $J_{i}$.
-Consider a three-input ($I=3$) and one output ($L=1$) system and assume the input domain is discretized with $J=J_{1}\times J_{2}\times J_{3}$ nodes, as shown in the special case box of Figure 1.
-To facilitate tensor notation, the nodal values will be denoted with left/right super/sub scripts, $\prescript{m}{i}{u}^{(j)}_{l}$, where $i\in\mathbb{N}^{I}$ is the input index, $m\in\mathbb{N}^{M_{i}}$ is the mode index, $l\in\mathbb{N}^{L}$ is the output index, and $j\in\mathbb{N}^{J_{i}}$ is the nodal index.
-The interpolated field $\mathcal{J}u(\bm{x})\in\mathbb{R}^{L=1}$ can be represented as a Tucker product:
-$\mathcal{J}u(\bm{x})=\llbracket\bm{\mathcal{G}};(\mathcal{J}\prescript{}{1}{\bm{u}})^{T},(\mathcal{J}\prescript{}{2}{\bm{u}})^{T},(\mathcal{J}\prescript{}{3}{\bm{u}})^{T}\rrbracket=\bm{\mathcal{G}}\times^{2}_{1}(\mathcal{J}\prescript{}{1}{\bm{u}})^{T}\times^{2}_{2}(\mathcal{J}\prescript{}{2}{\bm{u}})^{T}\times^{2}_{3}(\mathcal{J}\prescript{}{3}{\bm{u}})^{T},$ (4)
-where the core tensor $\bm{\mathcal{G}}\in\mathbb{R}^{M_{1}\times M_{2}\times M_{3}}$ is a trainable full matrix typically smaller than the original tensor that compresses data.
-The $\mathcal{J}\prescript{}{i}{\bm{u}}(x_{i})\in\mathbb{R}^{M_{i}\times 1}$ is one-dimensional (1D) interpolated output of $i$-th input dimension over $M_{i}$ modes represented as:
-$\mathcal{J}\prescript{}{i}{\bm{u}}(x_{i})=\begin{bmatrix}\prescript{1}{i}{u}^{(1)}&\prescript{1}{i}{u}^{(2)}&\cdots&\prescript{1}{i}{u}^{(J_{i})}\\ \prescript{2}{i}{u}^{(1)}&\prescript{2}{i}{u}^{(2)}&\cdots&\prescript{2}{i}{u}^{(J_{i})}\\ \vdots&\vdots&\vdots&\vdots\\ \prescript{M_{i}}{i}{u}^{(1)}&\prescript{M_{i}}{i}{u}^{(2)}&\cdots&\prescript{M_{i}}{i}{u}^{(J_{i})}\end{bmatrix}\cdot\begin{bmatrix}N^{(1)}(x_{i})\\ N^{(2)}(x_{i})\\ \vdots\\ N^{(J_{i})}(x_{i})\end{bmatrix}=\prescript{}{i}{\bm{U}}\times^{1}_{2}\prescript{}{i}{\bm{\mathcal{X}(x)}}=\prescript{}{i}{\bm{U}}\prescript{}{i}{\bm{\mathcal{X}(x)}},$ (5)
-where $\prescript{}{i}{\bm{U}}\in\mathbb{R}^{M_{i}\times J_{i}}$ and $\prescript{}{i}{\bm{\mathcal{X}(x)}}\in\mathbb{R}^{J_{i}}$.
-As illustrated in Figure 1’s special case box, the message passing (blue arrow) only happens in the axial directions, yielding 1D interpolation functions: $N^{(J_{i})}(x_{i})$.
-When there are more than one outputs $(L>1)$, the interpolated field becomes a vector of $L$ elements:
-$\mathcal{J}\bm{u}(\bm{x})$ $=\left[\mathcal{J}u_{1}(\bm{x}),\mathcal{J}u_{2}(\bm{x}),\cdots\mathcal{J}u_{L}(\bm{x})\right],\textit{where}$ (6)
-$\mathcal{J}u_{l}(\bm{x})$ $=\llbracket\bm{\mathcal{G}}_{l};(\mathcal{J}\prescript{}{1}{\bm{u}}_{l})^{T},(\mathcal{J}\prescript{}{2}{\bm{u}}_{l})^{T},(\mathcal{J}\prescript{}{3}{\bm{u}}_{l})^{T}\rrbracket=\bm{\mathcal{G}}_{l}\times^{2}_{1}(\mathcal{J}\prescript{}{1}{\bm{u}}_{l})^{T}\times^{2}_{2}(\mathcal{J}\prescript{}{2}{\bm{u}}_{l})^{T}\times^{2}_{3}(\mathcal{J}\prescript{}{3}{\bm{u}}_{l})^{T},$
-where $\mathcal{J}\prescript{}{i}{\bm{u}}_{l}\in\mathbb{R}^{M_{i}\times 1}$.
-The trainable parameters are the core tensors $\mathcal{G}_{l}\in\mathbb{R}^{M_{1}\times M_{2}\times M_{3}},l=1,\dots,L$, and the nodal values $\prescript{}{i}{\bm{U}}_{l}\in\mathbb{R}^{M_{i}\times J_{i}},l=1,\dots,L$, yielding a total count of $L\left(\prod^{I}_{i}{M_{i}}+\sum^{I}_{i}{M_{i}J_{i}}\right)$ that scales linearly with the nodal discretization $J_{i}$.
-Special case 2: regular mesh with CP decomposition
-Tucker decomposition can be further simplified to CANDECOMP/PARAFAC (CP) decomposition [38, 39, 40] by setting the core tensor $\bm{\mathcal{G}}$ as an order-$I$ super diagonal tensor: $\bm{\mathcal{G}}\in\mathbb{R}^{M^{I}}$, $M=M_{1}=\cdots=M_{I}$, all zero entries except the diagonal elements.
-If we further set the diagonal elements of $\bm{\mathcal{G}}$ to be 1, the Tucker decomposition in Eq.
-6 becomes:
-$\mathcal{J}\bm{u}(\bm{x})=\sum^{M}_{m=1}{[\mathcal{J}\prescript{m}{1}{\bm{u}}(x_{1})\odot\dots\odot\mathcal{J}\prescript{m}{I}{\bm{u}}(x_{I})]},$ (7)
-where $\mathcal{J}\prescript{m}{i}{\bm{u}}(x_{i})\in\mathbb{R}^{L}$ and $\odot$ represents element-wise multiplication.
-In CP decomposition, the core tensor is no longer trainable, thus the total trainable parameter becomes $ML\sum^{I}_{i}{J_{i}}$.
-It is important to note that the interpolation function $N^{(j)}(x_{i})$ in Eq.5 is one-dimensional.
-In other words, both Tucker and CP decomposition replace a high-dimensional interpolation with one-dimensional interpolations that dramatically reduce the number of trainable parameters (or DoFs).
-As given in Table 1, the number of trainable parameters of the full interpolation scales exponentially with the discretization $J_{i}$, whereas the Tucker and the CP decomposition scale linearly with $J_{i}$.
-Considering the trainable parameters of multi-layer perception (MLP) scales quadratically with the number of hidden neurons, INNs with tensor decomposition (TD) may dramatically reduce the model complexity and computing requirements.
-Table 1: Number of trainable parameters (or degrees of freedom, DoFs) of full interpolation, Tucker decomposition, and CP decomposition for an $I$-input single output $(L=1)$ relationship.
-The input domain is discretized with a regular grid of $J_{1}\times J_{2}\times\cdots\times J_{I}$ discretization.
-We assume that the nodal coordinates $(\bm{x}^{(j)})$ are fixed.
-Full interpolation Tucker decomposition CP decomposition
-Trainable parameters $\prod^{I}_{i}{J_{i}}$ $\prod^{I}_{i}{M_{i}}+\sum^{I}_{i}{M_{i}J_{i}}$ $M\sum^{I}_{i}{J_{i}}$
-INN trainer
-A general INN forward propagation is provided in Eq.3 as a tensor contraction.
-The INN trainer optimizes nodal values $\bm{U}=\{\bm{u}^{(j)}\}_{j=1,\cdots,J}$ (and if needed, nodal coordinates ($\bm{X}=\{\bm{x}^{(j)}\}_{j=1,\cdots,J}$)) under a given loss function and training data.
-Consider a regression problem with $K$ labeled data: $(\bm{x}^{*}_{k},\bm{u}^{*}_{k})$, $k=1,\cdots,K$; and $\bm{x}^{*}_{k}\in\mathbb{R}^{I},\bm{u}^{*}_{k}\in\mathbb{R}^{L}$.
-The superscript * denotes data.
-A mean squared error (MSE) loss function for this regression problem is defined as:
-$loss(\bm{U},\bm{X})=\frac{1}{K}\sum_{k}(\mathcal{J}\bm{u}(\bm{x}^{*}_{k})-\bm{u}^{*}_{k})^{2}.$ (8)
-Finally, an optimization is formulated as:
-$\underset{\bm{U},\bm{X}}{\text{minimize}}\ loss(\bm{U},\bm{X}).$ (9)
-Section 3.1 exhibits the application of INN trainer to metal additive manufacturing while other benchmarks are elucidated in SI Section 2.
-INN solver
-An INN solver generalizes classical numerical methods such as FEM and meshfree, as well as the model order reduction methods such as proper generalized decomposition (PGD) [41] and tensor decomposition (TD) [14, 42].
-Here, we introduce a formulation with CP decomposition to solve the generalized space $(\bm{x})$ time $(t)$ parameter $(\bm{\theta})$ (S-T-P) PDE, whose computational cost is prohibitively high for most numerical methods and machine learning approaches [41].
-Consider a parameterized space-time PDE:
-$\mathcal{L}\bm{u}(\bm{x},t,\bm{\theta})=0,$ (10)
-where $\mathcal{L}$ is the general partial differential operator (can be linear or nonlinear).
-The INN solver uses the same neural network structure as the trainer, but the loss function varies with the equations to be solved.
-It is formulated as the weighted summation of the PDE residual [43, 44, 45].
-An INN solution field in the S-T-P domain can be written as:
-$\begin{split}\mathcal{J}{\bm{u}}(\bm{x},t,\bm{\theta};\bm{U},\bm{X})&=\sum_{m=1}^{M}[\mathcal{J}\prescript{m}{x_{1}}{\bm{u}}(x_{1})]\odot\cdots\odot[\mathcal{J}\prescript{m}{x_{d}}{\bm{u}}(x_{d})]\odot[\mathcal{J}\prescript{m}{t}{\bm{u}}(x_{t})]\\ &\odot[\mathcal{J}\prescript{m}{\theta_{1}}{\bm{u}}(\theta_{1})]\odot\cdots\odot[\mathcal{J}\prescript{m}{\theta_{k}}{\bm{u}}(\theta_{k})],\end{split}$ (11)
-where $d$ and $k$ are the spatial dimension and the number of parameters, respectively.
-Similar to the trainer, the goal is to find nodal values $\bm{U}$ (and nodal coordinates $\bm{X}$, if one wants to adapt the mesh).
-As a result, the INN obtains the S-T-P solution by minimizing the loss function:
-[…]
-where $\delta\bm{u}$ is the weight function and can be defined using Galerkin, Petrov-Galerkin, collocation or quadratic formulation.
-Due to the Kronecker delta property of INN interpolation functions, Dirichlet boundary conditions and initial conditions can be strongly imposed.
-As a result, only the weighted sum residual is considered in Eq.
-11 as a loss function.
-This distinguishes INNs from most data-driven PDE solvers that weakly impose those conditions [46].
-See the SI Section 4 for detailed derivations of the loss function and solution scheme.
-INN solvers applied to the metal additive manufacturing and the linear elastic solid mechanics problem are introduced in Section 3.3 and SI Section 4, respectively.
-INN calibrator
-[…]
-$\bm{x}^{*}=\underset{\bm{x}}{\text{argmin}}\frac{1}{K}\sum_{k}(\mathcal{J}\bm{u}(\bm{x})-\ \bm{u}^{*}_{k})^{2},$ (13)
-where $\mathcal{J}\bm{u}(\bm{x})$ is the trained or solved forward model, $\bm{u}^{*}_{k}$ is the $k$-th measured observation and $\bm{x}^{*}$ is the calibrated input variable.
-The INN calibrator applied to the heat source calibration task in metal additive manufacturing can be found in Section 3.3 while other benchmarks can be found in SI Section 3.
-3 Scientific and Engineering applications: Metal Additive Manufacturing
+Once the forward propagation is defined, the loss function is chosen based on the problem type: training, solving, or calibrating (see Section 5 for details).
+When the input domain is discretized with a regular mesh, as illustrated in the special case of Figure 1, we can significantly reduce the trainable parameters (or degrees of freedom, DoFs) by leveraging Tensor Decomposition (TD) [19].
+Section 5 introduces the two widely accepted TD methods: Tucker decomposition [33, 34] and CANDECOMP/PARAFAC (CP) decomposition [35, 36, 37].
+Figure 3: Benchmarks for the INN solver (a) and the INN trainer (b).
+In (a), a 3D-space, time, and 4D-parameter heat transfer equation is solved to model a 10 mm single-track laser powder bed fusion (L-PBF) metal AM.
+Detailed problem definition and explanation can be found in Section 3.
+We compare the single-scale finite difference method (FDM) solver [24], the variational multiscale FEM solver [11], and the INN solver with CP decomposition and Q=2.
+The data points for the first two methods with dashed marker edges are estimated, while that of INN with a solid marker edge is computed.
+In (b), a standard data training (regression) problem is studied.
+Each trainer stops when the training loss hits the stopping criteria (MSE:4e-4).
+We set the same optimization condition: ADAM optimizer; learning rate: 1e-3; batch size: 128.
+The number of neurons per hidden layer is denoted for each MLP data point while the number of segments and modes are denoted for each INN data point.
+INN with CP decomposition and Q=2 is used.
+In (c), a 1D Poisson’s equation is solved with PINN and INN.
+PINN is made of a 1-layer MLP with a varying number of neurons.
+Randomly selected 10k collocation points are used to compute the PDE loss.
+With a batch size of 128, 10k epochs are trained using an ADAM optimizer with a learning rate of 1e-1.
+FEM with linear elements is equivalent to an INN with Q=1, s=1.
+C-HiDeNNs with s=3,P=3 and with s=5,P=5 are equivalent to INNs with Q=2, s=3, P=3, and with Q=2, s=5, P=5, respectively.
+All solvers and trainers are graphics processing units (GPU) optimized with the JAX library [38].
+2.1 Benchmarking INN trainer
+To highlight INN’s advantages in speed (training epoch) and storage (number of parameters) over traditional MLP, we introduce a benchmark problem: a 10-input 5-output physical function [39], (see SI Section 2.4 for details).
+Using this deterministic equation, we randomly generate 100,000 data using a Latin hypercube sampling.
+The dataset is split into 80% and 20% for train and test, respectively.
+MLPs with two and three hidden layers and with a sigmoid activation are tested while INNs with CP decomposition, two (Q=2) message passing, and $s=2,P=2$ polynomial activation are adopted.
+To investigate the convergence behavior depending on the number of parameters, we set the stopping criteria as training MSE: 4e-4 and count the epoch at convergence.
+Since the physical function is deterministic (no noise) and we drew a sufficiently large number of data, no considerable overfitting was observed.
+Figure 3(b) reveals that given the same number of trainable parameters, INNs converge significantly faster than MLPs.
+For instance, the INN with 2,250 parameters converges at the 34th epoch while the MLP with 2,285 parameters converges at the 217th epoch.
+MLPs with more parameters tend to stop at earlier epochs, however, even the largest MLP with 21,805 parameters converges at the 45th epoch while that of INN with only 8,250 parameters converges at the 13th epoch.
+This benchmark demonstrates that the INN trainer is lightweight and fast-converging compared to traditional MLPs.
+Other benchmarks of the INN trainer are elucidated in SI Section 2.
+2.2 Benchmarking INN solver
+INNs can achieve a proven convergence rate when solving a PDE.
+In numerical analysis, the convergence rate is the decaying rate of an error measurement as the mesh refines.
+This benchmark solves a 1D Poisson’s equation defined in Section 5.6 with the $H^{1}$ norm error estimator.
+It is mathematically proven that a numerical solution of FEM or any interpolation-based solution method exhibits a convergence rate of $P$, which is the complete order of polynomial basis used in the interpolation [21].
+Figure 3(c) is the $H^{1}$ error vs.
+parameters (i.e., degrees of freedom) plot where the convergence rates are denoted as the slope of the log-log plot.
+Depending on the message passing operation, INNs degenerate to linear FEM ($P=1$) or C-HiDeNNs (or GFEM) with different order of polynomials: $P=3,P=5$.
+However, PINNs constructed with MLP do not reveal a convergent behavior.
+Only a few theoretical works have proved the convergence rate of PINNs under specific conditions [40].
+A PDE solver needs to have a stable and fast convergence rate because it guides an engineer in choosing the mesh resolution and other hyperparameters for achieving the desired level of accuracy.
+We also demonstrate the convergent behavior of INN solvers with and without CP decomposition for a 3D linear elasticity equation in SI Sections 4.5 and 4.6.
+3 Discussion
 Figure 4: A summary of engineering applications of INNs to additive manufacturing (AM) problems.
@@
 Finally, the INN solver solves a space-time-parameter heat transfer equation, resulting in a significant storage reduction and faster simulations.
-This section illustrates how Engineering Software 2.0 enabled with INNs can be applied to advance various facets of metal additive manufacturing (AM).
+Now we illustrate how Engineering Software 2.0 enabled with INNs can be applied to advance various facets of metal additive manufacturing (AM).
 Note that the INN with CP decomposition will be referred to as INN in this section for brevity.
-Laser powder bed fusion (L-PBF) is a type of metal additive manufacturing in which metal powders are layered and melted by a laser beam to create objects that match the provided CAD designs [47].
+Laser powder bed fusion (L-PBF) is a type of metal AM in which metal powders are layered and melted by a laser beam to create objects that match the provided CAD designs [41].
 The added flexibility of L-PBF comes with a huge design space and spatially varying microstructure after manufacturing.
@@
 The goal is to maintain a homogeneous temperature of the melt pool (i.e., a molten region near the laser spot) across each layer using real-time sensor data, under the hypothesis that a homogeneous melt pool temperature across each layer of material deposition will reduce the variability in the microstructure and mechanical properties.
-The laser power at each point of a certain layer needs to be controlled (increased/decreased by certain amount compared to previous layer) to achieve this homogeneity in melt pool temperature.
-The fundamental challenge of implementing a model predictive control system for such an application is the computational resources required for forward prediction (involving finite element or computational fluid dynamics models) and inverse prediction, as illustrated in Figure 3.
+The laser power at each point of a certain layer needs to be controlled (increased/decreased by a certain amount compared to the previous layer) to achieve this homogeneity in melt pool temperature.
+The fundamental challenge of implementing a model predictive control system for such an application is the computational resources required for forward prediction (involving finite element or computational fluid dynamics models) and inverse prediction, as illustrated in Figure 4, the row for training.
 The INN trainer aims to provide a reliable and memory-efficient surrogate model that can replace mechanistic models in predictive control systems.
@@
 $S_{h}(\mathbf{x},t,\mathbf{p})=\frac{2P\eta}{\pi r^{2}_{b}d}e^{-\frac{2(x^{2}+y^{2})}{r^{2}_{b}}}\cdot\mathbf{1}_{\{z|z_{top}-z\leq d\}}(z)$ (5)
-where $z_{top}$ is the z-coordinate of the current surface of the powder bed and $\mathbf{1}$ is the indicator function.
+where $z_{top}$ is the z-coordinate of the current surface of the powder bed and $\mathbf{1}$ is the indicator function that returns 1 if the condition is met and 0 otherwise.
 The notation follows Eq.
-14, and $r_{b}$ is the Gaussian profile’s standard deviation that controls the beam’s width.
+4, and $r_{b}$ is the Gaussian profile’s standard deviation that controls the beam’s width.
 Following reference [10], the heat source parameters $d,\eta,r_{b}$ are controlled by the calibration parameters $p_{1}$, $p_{2}$, $p_{3}$ such that $d=p_{1}\frac{P}{V}RHF^{2}$, $\eta=p_{2}\frac{P}{V}RHF^{2}$, and $r_{b}=p_{3}\frac{P}{V}RHF^{2}$.
 Here, $V$ is the laser scan speed and $RHF$ is a predetermined residual heat factor based on the laser toolpath [45].
-This calibration problem’s major challenge is achieving a highly accurate surrogate model given a sparse set of training data.
-INNs outperform MLPs in solving these data-scarce problems because they are rooted in the interpolation theory.
+The major challenge of this problem is to achieve a highly accurate surrogate model given a sparse set of training data.
+INNs outperform MLPs in training these scarce data because they are rooted in the interpolation theory.
 As a demonstration, an INN with 1760 trainable parameters is trained on 28 sets of meltpool temperature data generated by the finite difference method (FDM) [24] using 4 different power-scanning speed set-ups.
 Later, this trained INN is used as a surrogate to optimize the three parameters ($p_{1},p_{2},p_{3}$) given the experimental data.
-To show the superiority of INNs over traditional neural networks in such data-scarce applications, a 3-layer MLP with 2689 trainable parameters (roughly the same order of magnitude as the INN) is trained using the same data fed to the INN.
-In practice, we observed that INN converges well within 100 epoch, while MLP requires over 1000 epoch to converge to a test MSE 10 times larger than INN given all other setup the same.
+For comparison, a 3-layer MLP with 2689 trainable parameters (roughly the same order of magnitude as the INN) is trained using the same data.
+We observed that INNs converge well within 100 epochs, while MLPs require over 1000 epochs to converge to a test MSE 10 times larger than the INN given all other setups the same.
 Details on this training and calibration process are provided in SI Section 3.1.
-The INN- and MLP-based calibrations are repeated 50 times to provide a statistical comparison (see Figure 3).
+The INN- (definition of the INN calibrator can be found in Section 5.5 and MLP-based calibrations are repeated 50 times to provide a statistical comparison (see Figure 4).
 Even with only 28 sets of simulation data, the calibrated parameters using the INN produce a mean melt pool temperature within a 7.9% difference of the experimental data.
@@
 On the other hand, the INN achieves a much lower difference, showing that the model can learn underlying physics even with a scarce dataset.
-3.3 Solving High Dimensional Space-Time-Parameter Heat Equation
+3.3 Solving high dimensional Space-Time-Parameter (S-T-P) heat equation
 […]
-The surrogate model $f$ can also be obtained using the classical data-driven approach.
-In the offline stage, training data can be generated by repeatedly running numerical solvers (such as finite difference method (FDM) and variational multiscale method (VMS)) in the Space-Time domain with different sets of parameters $\mathbf{{p}}$.
+The surrogate model $f$ can also be obtained using the standard data-driven approach.
+In the offline stage, training data are generated by repeatedly running numerical solvers (such as FDM or variational multiscale method (VMS)) in the Space-Time domain with different sets of parameters $\mathbf{{p}}$.
 Then a surrogate model is trained with the data.
-The standard data-driven approach suffers from the curse of dimensionality when the parametric space is high-dimensional.
-As a result, the data-driven approaches face three challenges at the offline stage: expensive computation for repetitive data generation, huge memory cost for running full-scale simulations, and excessive disk storage of offline training data.
-We estimate the time and storage required for the data-driven approach with FDM and VMS, as shown in Table 2 and illustrated in Figure 2.
+The data-driven approach suffers from the curse of dimensionality when the parametric space is high-dimensional, resulting in expensive computation for repetitive data generation, huge memory costs for running full-scale simulations, and excessive disk storage of offline training data.
+We estimate the time and storage required for the data-driven approaches with FDM and VMS, as shown in Table 1 and illustrated in Figure 3(a).
 Unlike the data-driven approaches, INN solvers treat the parameters $\mathbf{p}$ as additional parametric inputs.
-As a result, INNs can directly obtain a parametric surrogate model $f=u(\mathbf{x},t,\mathbf{p})$ directly from the governing equation without going through the cumbersome offline data generation and surrogate model training process.
-This is particularly beneficial to high-dimensional parametric surrogate models.
+As a result, INNs obtain a parametric surrogate model $f=u(\mathbf{x},t,\mathbf{p})$ directly from the governing equation without going through the cumbersome offline data generation and surrogate model training.
 Moreover, our results prove the INN solution (or the S-T-P surrogate model) achieves an exceptionally high accuracy of $R^{2}$ = 0.9969, considering that most data-driven approaches for S-T-P problems trained on numerical simulation data suffer from low model accuracy, typically below $R^{2}=0.9$ [46].
@@
 The performance comparison of INN versus data-driven methods is shown in Table.
-2.
+1.
 Table 1: Performance comparison of different surrogate models.
@@
 INN 8 $1$ 15 minutes $7.37\times 10^{-2}$
-4 Conclusions
-This article demonstrates that the interpolating neural network (INN) can train, solve, and calibrate scientific and engineering problems that are extremely challenging or prohibitive for existing numerical methods and machine learning models.
-The keys to INN’s success are 1) it interpolates the graph nodes’ values with well-established interpolation theories (not interpolating training data with a black-box network that often causes overfitting issues) and 2) it leverages the tensor decomposition (TD) to resolve curse of dimensionality.
-Due to the significantly reduced trainable parameters without losing approximation capabilities, INN can train/solve/calibrate faster with less memory requirement, and achieve extremely high model accuracy and fast inference time.
-As this is the original article on INNs, we have myriads of research questions that may gather huge research interests across various areas including but not limited to scientific machine learning, applied mathematics, computer vision, and data science.
+4 Conclusion
+This article demonstrates that interpolating neural networks (INNs) can train, solve, and calibrate scientific and engineering problems that are extremely challenging or prohibitive for existing numerical methods and machine learning models.
+The keys to INN’s success are 1) it interpolates the graph nodes’ values with well-established interpolation theories and 2) it leverages the tensor decomposition (TD) to resolve the curse of dimensionality.
+Due to the reduced number of parameters without compromising predictive accuracy, INNs become an efficient substitute for MLPs or PINNs.
+There are numerous research questions that could generate significant interest across various fields, including, but not limited to, scientific machine learning, applied mathematics, computer vision, and data science.
 To mention a few:
-  • Depending on the problem, INN’s superior training efficiency may cause overfitting issues.
-The ensemble training (i.e., training on random subsets of the training data) might be a way to enhance the generality of the model for image problems or noisy data.
-  • Multi-resolution aspect (i.e., a varying number of segments across TD modes) of INNs needs to be investigated. See SI 1.4 for further discussions.
-  • Similar to the INN solvers, the convergence of INN trainers can be studied mathematically.
+  • Depending on the problem, INN’s superior training efficiency may overfit data.
+The ensemble training (i.e., training on random subsets of the training data) might be a way to enhance the generality of the INN model for real-world data.
+  • Multi-resolution aspect (i.e., a varying mesh resolution across TD modes) of INNs needs to be investigated. See SI 1.4 for further discussions.
+  • The interpretability of INNs should be investigated.
+INNs may address the challenge of model-based interpretability (a terminology defined from [47]), which involves developing models that are both simple enough to be easily understood by developers and capable of maintaining high predictive accuracy.
+  • Similar to INN solvers, the convergence of INN trainers can be studied both mathematically and numerically.
   • Since INNs can be used as both a solver for physics-based problems and a function approximator for data-driven problems, it is of interest to combine the two behaviors into one single model to solve large-scale problems involving both complex physics and scarce data.
-  • The superior performance of the INN solver may facilitate multiscale analysis in a vast parametric space.
+  • The superior performance of INN solvers may facilitate multiscale analysis in a vast parametric space.
 One can integrate the S-T-P INN solver with a concurrent multiscale analysis framework such as self-consistent clustering analysis [48, 49] and open a new direction towards parameterized multiscale analysis.
   • The current INN code has been optimized to run on a single GPU. Since the INN forward pass over multiple modes can be parallelized, multi-GPU programming will further speed up the code.
-  • Although Tucker decomposition has been introduced as a special case 1 in Section 2, the numerical experiments conducted in this article mostly focus on the CP decomposition.
+  • Although Tucker decomposition is introduced in Section 5.1, the numerical experiments conducted in this article mostly focus on the CP decomposition.
 Theoretically, Tucker decomposition has a larger approximation space than CP decomposition.
@@
 We expect INNs will open a next-generation Engineering Software 2.0 that revolutionizes almost all fields of computational science and engineering.
+5 Methods
+5.1 Special case 1: regular mesh with Tucker decomposition
+One of the widely used TD methods is Tucker decomposition [33, 34].
+It approximates a high-order tensor with $J_{i}$ nodes in $i$-th dimension as a tensor contraction between dimension-wise matrices and a core tensor $\bm{\mathcal{G}}$, which has the same order of the original tensor but with smaller nodes $M_{i}$ ($<J_{i}$).
+The $M_{i}$ is often called a "mode" to distinguish it from the original node $J_{i}$.
+Consider a three-input ($I=3$) and one output ($L=1$) system, and assume the input domain is discretized with $J=J_{1}\times J_{2}\times J_{3}$ nodes, as shown in the special case box of Figure 1.
+To facilitate tensor notation, the nodal values will be denoted with left/right super/sub scripts, $\prescript{m}{i}{u}^{(j)}_{l}$, where $i\in\mathbb{N}^{I}$ is the input index, $m\in\mathbb{N}^{M_{i}}$ is the mode index, $l\in\mathbb{N}^{L}$ is the output index, and $j\in\mathbb{N}^{J_{i}}$ is the nodal index.
+The interpolated field $\mathcal{J}u(\bm{x})\in\mathbb{R}^{L=1}$ can be represented as a Tucker product:
+$\mathcal{J}u(\bm{x})=\llbracket\bm{\mathcal{G}};(\mathcal{J}\prescript{}{1}{\bm{u}})^{T},(\mathcal{J}\prescript{}{2}{\bm{u}})^{T},(\mathcal{J}\prescript{}{3}{\bm{u}})^{T}\rrbracket=\bm{\mathcal{G}}\times^{2}_{1}(\mathcal{J}\prescript{}{1}{\bm{u}})^{T}\times^{2}_{2}(\mathcal{J}\prescript{}{2}{\bm{u}})^{T}\times^{2}_{3}(\mathcal{J}\prescript{}{3}{\bm{u}})^{T},$ (6)
+where the core tensor $\bm{\mathcal{G}}\in\mathbb{R}^{M_{1}\times M_{2}\times M_{3}}$ is a trainable full matrix typically smaller than the original tensor that compresses the data.
+Here, $\bm{\mathcal{A}}\times^{b}_{a}\bm{\mathcal{B}}$ denotes the tensor contraction operation between the $a$-th dimension of tensor $\bm{\mathcal{A}}$ and the $b$-th dimension of tensor $\bm{\mathcal{B}}$.
+The $\mathcal{J}\prescript{}{i}{\bm{u}}(x_{i})\in\mathbb{R}^{M_{i}\times 1}$ is one-dimensional (1D) interpolated output of $i$-th input dimension over $M_{i}$ modes represented as:
+$\mathcal{J}\prescript{}{i}{\bm{u}}(x_{i})=\begin{bmatrix}\prescript{1}{i}{u}^{(1)}&\prescript{1}{i}{u}^{(2)}&\cdots&\prescript{1}{i}{u}^{(J_{i})}\\ \prescript{2}{i}{u}^{(1)}&\prescript{2}{i}{u}^{(2)}&\cdots&\prescript{2}{i}{u}^{(J_{i})}\\ \vdots&\vdots&\vdots&\vdots\\ \prescript{M_{i}}{i}{u}^{(1)}&\prescript{M_{i}}{i}{u}^{(2)}&\cdots&\prescript{M_{i}}{i}{u}^{(J_{i})}\end{bmatrix}\cdot\begin{bmatrix}N^{(1)}(x_{i})\\ N^{(2)}(x_{i})\\ \vdots\\ N^{(J_{i})}(x_{i})\end{bmatrix}=\prescript{}{i}{\bm{U}}\times^{1}_{2}\prescript{}{i}{\mathcal{X}(x_{i})}=\prescript{}{i}{\bm{U}}\prescript{}{i}{\mathcal{X}(x_{i})},$ (7)
+where $\prescript{}{i}{\bm{U}}\in\mathbb{R}^{M_{i}\times J_{i}}$ and $\prescript{}{i}{\bm{\mathcal{X}(x)}}\in\mathbb{R}^{J_{i}}$.
+As illustrated in the special case box of Figure 1, the message passing (blue arrow) only happens in the axial directions, yielding 1D interpolation functions: $N^{(J_{i})}(x_{i})$.
+When there are more than one output $(L>1)$, the interpolated field becomes a vector of $L$ elements:
+$\mathcal{J}\bm{u}(\bm{x})$ $=\left[\mathcal{J}u_{1}(\bm{x}),\mathcal{J}u_{2}(\bm{x}),\cdots\mathcal{J}u_{L}(\bm{x})\right],\textit{where}$ (8)
+$\mathcal{J}u_{l}(\bm{x})$ $=\llbracket\bm{\mathcal{G}}_{l};(\mathcal{J}\prescript{}{1}{\bm{u}}_{l})^{T},(\mathcal{J}\prescript{}{2}{\bm{u}}_{l})^{T},(\mathcal{J}\prescript{}{3}{\bm{u}}_{l})^{T}\rrbracket=\bm{\mathcal{G}}_{l}\times^{2}_{1}(\mathcal{J}\prescript{}{1}{\bm{u}}_{l})^{T}\times^{2}_{2}(\mathcal{J}\prescript{}{2}{\bm{u}}_{l})^{T}\times^{2}_{3}(\mathcal{J}\prescript{}{3}{\bm{u}}_{l})^{T},$
+where $\mathcal{J}\prescript{}{i}{\bm{u}}_{l}\in\mathbb{R}^{M_{i}\times 1}$.
+The trainable parameters are the core tensors $\mathcal{G}_{l}\in\mathbb{R}^{M_{1}\times M_{2}\times M_{3}},l=1,\dots,L$, and the nodal values $\prescript{}{i}{\bm{U}}_{l}\in\mathbb{R}^{M_{i}\times J_{i}},l=1,\dots,L$, yielding a total count of $L\left(\prod^{I}_{i}{M_{i}}+\sum^{I}_{i}{M_{i}J_{i}}\right)$ that scales linearly with the nodal discretization $J_{i}$.
+5.2 Special case 2: regular mesh with CP decomposition
+Tucker decomposition can be further simplified to CANDECOMP/PARAFAC (CP) decomposition [35, 36, 37] by setting the core tensor $\bm{\mathcal{G}}$ as an order-$I$ super diagonal tensor: $\bm{\mathcal{G}}\in\mathbb{R}^{M^{I}}$, $M=M_{1}=\cdots=M_{I}$, all zero entries except the diagonal elements.
+If we further set the diagonal elements of $\bm{\mathcal{G}}$ to be 1, the Tucker decomposition in Eq.
+8 becomes:
+$\mathcal{J}\bm{u}(\bm{x})=\sum^{M}_{m=1}{[\mathcal{J}\prescript{m}{1}{\bm{u}}(x_{1})\odot\dots\odot\mathcal{J}\prescript{m}{I}{\bm{u}}(x_{I})]},$ (9)
+where $\mathcal{J}\prescript{m}{i}{\bm{u}}(x_{i})\in\mathbb{R}^{L}$ and $\odot$ represent multiplication in elements.
+In CP decomposition, the core tensor is no longer trainable, thus the total trainable parameter becomes $ML\sum^{I}_{i}{J_{i}}$.
+It is important to note that the interpolation function $N^{(j)}(x_{i})$ in Eq.7 is one-dimensional.
+In other words, both Tucker and CP decomposition replace a high-dimensional interpolation with one-dimensional interpolations that dramatically reduce the number of trainable parameters (or DoFs).
+As given in Table 2, the number of trainable parameters of the full interpolation scales exponentially with the discretization $J_{i}$, whereas the Tucker and CP decomposition scale linearly with $J_{i}$.
+Considering the trainable parameters of multi-layer perceptron (MLP) scales quadratically with the number of hidden neurons, INNs with tensor decomposition (TD) may dramatically reduce the model complexity and computing requirements.
+However, the reconstructed basis using TD loses expressibility because of the separated variables.
+There is no cross-term in the functional space of TD.
+Nevertheless, we can transform this into an advantage by enriching the 1D approximations with adaptive activation (see SI Section 2.2) for the Q-step message passing.
+Therefore, we can find a hybrid architecture that preserves some of the advantages of numerical methods (i.e., FEM) and neural network methods while acknowledging that it loses something from both.
+Table 2: Number of trainable parameters (or degrees of freedom, DoFs) of full interpolation, Tucker decomposition, and CP decomposition for an $I$-input single output $(L=1)$ relationship.
+The input domain is discretized with a regular grid of $J_{1}\times J_{2}\times\cdots\times J_{I}$ discretization.
+We assume that the nodal coordinates $(\bm{x}^{(j)})$ are fixed.
+Full interpolation Tucker decomposition CP decomposition
+Trainable parameters $\prod^{I}_{i}{J_{i}}$ $\prod^{I}_{i}{M_{i}}+\sum^{I}_{i}{M_{i}J_{i}}$ $M\sum^{I}_{i}{J_{i}}$
+5.3 Loss function of INN trainer
+A general INN forward propagation is provided in Eq.3 as a tensor contraction.
+INN can be used for data training as any other neural network architecture.
+While MLPs optimize weights and biases during training, INN trainers optimize nodal values $\bm{U}=\{\bm{u}^{(j)}\}_{j=1,\cdots,J}$ (and, if needed, nodal coordinates ($\bm{X}=\{\bm{x}^{(j)}\}_{j=1,\cdots,J}$)) under a given loss function and training data.
+Consider a regression problem with $K$ labeled data: $(\bm{x}^{*}_{k},\bm{u}^{*}_{k})$, $k=1,\cdots,K$; and $\bm{x}^{*}_{k}\in\mathbb{R}^{I},\bm{u}^{*}_{k}\in\mathbb{R}^{L}$.
+The superscript * denotes the data.
+A mean squared error (MSE) loss function for this regression problem is defined as:
+$loss(\bm{U},\bm{X})=\frac{1}{K}\sum_{k}(\mathcal{J}\bm{u}(\bm{x}^{*}_{k})-\bm{u}^{*}_{k})^{2}.$ (10)
+Finally, an optimization is formulated as follows.
+$\underset{\bm{U},\bm{X}}{\text{minimize}}\ loss(\bm{U},\bm{X}).$ (11)
+5.4 Loss function of INN solver
+An INN solver generalizes classical numerical methods such as FEM and meshfree, as well as model order reduction methods such as proper generalized decomposition (PGD) [50] and tensor decomposition (TD) [20, 51].
+Here, we introduce a formulation with CP decomposition to solve the generalized space $(\bm{x})$ time $(t)$ parameter $(\bm{\theta})$ (S-T-P) PDE, whose computational cost is prohibitively high for most numerical methods and machine learning approaches [50].
+Consider a parameterized space-time PDE:
+$\mathcal{L}\bm{u}(\bm{x},t,\bm{\theta})=0,$ (12)
+where $\mathcal{L}$ is the general partial differential operator (can be linear or nonlinear).
+The INN solver uses the same neural network structure as the trainer, but the loss function varies with the equations to be solved.
+It is formulated as the weighted summation of the PDE residual [52, 53, 54].
+An INN solution field in the S-T-P domain can be written as:
+$\begin{split}\mathcal{J}{\bm{u}}(\bm{x},t,\bm{\theta};\bm{U},\bm{X})&=\sum_{m=1}^{M}[\mathcal{J}\prescript{m}{x_{1}}{\bm{u}}(x_{1})]\odot\cdots\odot[\mathcal{J}\prescript{m}{x_{d}}{\bm{u}}(x_{d})]\odot[\mathcal{J}\prescript{m}{t}{\bm{u}}(x_{t})]\\ &\odot[\mathcal{J}\prescript{m}{\theta_{1}}{\bm{u}}(\theta_{1})]\odot\cdots\odot[\mathcal{J}\prescript{m}{\theta_{k}}{\bm{u}}(\theta_{k})],\end{split}$ (13)
+where $d$ and $k$ are the spatial dimension and the number of parameters, respectively.
+Similar to the trainer, the goal is to find nodal values $\bm{U}$ (and nodal coordinates $\bm{X}$, if one wants to adapt the mesh).
+As a result, the INN obtains the S-T-P solution by minimizing the loss function:
+[…]
+where $\delta\bm{u}$ is the weight function and can be defined using Galerkin, Petrov-Galerkin, collocation or quadratic formulation.
+Due to the Kronecker delta property of INN interpolation functions, Dirichlet boundary conditions and initial conditions can be strongly imposed.
+As a result, only the weighted sum residual is considered in Eq.
+13 as a loss function.
+This distinguishes INNs from most data-driven PDE solvers that weakly impose these conditions [55].
+See Section 4 of the SI for detailed derivations of the loss function and solution scheme.
+INN solvers applied to metal additive manufacturing and the linear elastic solid mechanics problem are introduced in Section 3.3 and SI Section 4, respectively.
+5.5 Loss function of INN calibrator
+[…]
+$\bm{x}^{*}=\underset{\bm{x}}{\text{argmin}}\frac{1}{K}\sum_{k}(\mathcal{J}\bm{u}(\bm{x})-\ \bm{u}^{*}_{k})^{2},$ (15)
+where $\mathcal{J}\bm{u}(\bm{x})$ is the trained or solved forward model, $\bm{u}^{*}_{k}$ is the $k$-th measured observation and $\bm{x}^{*}$ is the calibrated input variable.
+The INN calibrator applied to the heat source calibration task in metal additive manufacturing can be found in Section 3.3 while other benchmarks can be found in SI Section 3.
+5.6 1D Poisson’s equation with error estimator
+We borrow the manufactured problem first introduced in [14]. The 1D Poisson’s equation is defined as:
+$\Delta u(x)+b(x)=0,\quad x\in[0,10],$ (16)
+$u(x=0)=u(x=10)=0,$
+where $\Delta$ is the Laplace operator. The manufactured solution becomes
+$u(x)=(e^{-\pi(x-2.5)^{2}}-e^{-6.25\pi})+2(e^{-\pi(x-7.5)^{2}}-e^{-56.25\pi})-\frac{e^{-6.25\pi}-e^{-56.25\pi}}{10}x$ (17)
+when the body force is given as:
+$b(x)=-\frac{4\pi^{2}(x-2.5)^{2}-2\pi}{e^{\pi(x-2.5)^{2}}}-\frac{8\pi^{2}(x-7.5)^{2}-4\pi}{e^{\pi(x-7.5)^{2}}}.$ (18)
+The $H^{1}$ norm error estimator is defined as:
+[…]
+where $u^{h}(x)$ is the approximated solution.
 Acknowledgments
@@
 Vispi Karkaria from Northwestern University for their invaluable comments and support throughout this work.
-We also acknowledge the following grants: DEVCOM Army Research Laboratory under Cooperative Agreement Numbers W911NF-20-2-0292 and W911NF-21-2-02199 to J.
-Cao and the startup funds provided by Virginia Polytechnic Institute and State University to S.
-Saha.
-C.
-Park acknowledges the internship opportunity with the Freudenberg Group, within the team managed by Dr.
-Vishwanath Hegadekatte.
+Code availability
+The computer codes used for the benchmarking can be found at https://github.com/hachanook/pyinn. Codes for additive manufacturing examples can be provided upon a reasonable request.